### v2

# Beyond Trend and Periodicity: Guiding Time Series Forecasting with Textual Cues

Abstract
This work introduces a novel Text-Guided Time Series Forecasting (TGTSF) task. By integrating textual cues, such as channel descriptions and dynamic news, TGTSF addresses the critical limitations of traditional methods that rely purely on historical data. To support this task, we propose TGForecaster, a robust baseline model that fuses textual cues and time series data using cross-attention mechanisms. We then present four meticulously curated benchmark datasets to validate the proposed task, ranging from simple periodic data to complex, event-driven fluctuations. Our comprehensive evaluations demonstrate that TGForecaster consistently achieves state-of-the-art performance, highlighting the transformative potential of incorporating textual information into time series forecasting. This work not only pioneers a novel forecasting task but also establishes a new benchmark for future research, driving advancements in multimodal data integration for time series models.
1 Introduction
Time series forecasting (TSF) is crucial in many fields, thereby attracting significant attention from both academia and industry. Despite extensive research dedicated to this task, recent studies have shown that simple linear models (Zeng et al., 2023; Xu et al., 2023; Toner & Darlow, 2024) that barely extract trend and periodicity information from time series data frequently achieve performance close to that of state-of-the-art complex models (Nie et al., 2023; Liu et al., 2023; Jin et al., 2023), indicating that current approaches have reached a saturation point, possibly overfitting historical data.
To achieve higher forecasting accuracy, it is essential to provide models with additional, critical information—such as holiday information, consumer sentiment, and climate changes (depending on the specific forecasting task)—that are not present in the time series data itself. Without this additional information, models cannot extract the necessary causal relationships, leading to oversimplification or overfitting. Moreover, the lack of insight about channels and underlying systems prevents the modeling of inter-channel relationships, often resulting in unreliable forecasts.
As such valuable information is usually hard to quantify as auxiliary time series data, we consider them as textual cues and introduce a novel task for TSF: Text-Guided Time Series Forecasting (TGTSF). TGTSF leverages two main components: channel descriptions and dynamic news messages. Channel descriptions provide static knowledge about the underlying systems, enabling the model to differentiate between channels and better understand inter-channel correlations. News messages offer dynamic and external insights, helping the model adjust to shifts in data distribution caused by external events. By modeling the joint distribution of future values conditioned on these textual cues, TGTSF has the potential to enhance forecasting accuracy and reliability. Furthermore, TGTSF enables scenario-based forecasting, which is particularly valuable for business decision-making. For instance, a company can forecast sales based on the assumption of a successful marketing campaign, incorporating relevant news and channel-specific descriptions to refine predictions.
We then introduce TGForecaster, a simple Transformer-based baseline multimodal model designed to realize the potential of the proposed TGTSF framework. Utilizing a cross-attention mechanism, TGForecaster fuses information from textual cues and time series data to guide the forecasting process. To rigorously evaluate and demonstrate the efficacy of our model, we carefully craft four benchmark datasets categorized into synthetic, captioned existing, and real-world datasets. Each dataset is uniquely designed to validate the task and test the model’s capability from different perspectives. Specifically, the synthetic dataset features simple yet variable time series patterns, such as sinusoidal waves with random frequency changes, to test the model’s ability to utilize textual cues in forecasting and mitigating distribution shifts. Two existing datasets are captioned with different levels of textual information, used to challenge the model in complex, non-periodic environments by integrating detailed external data. Lastly, our real-world dataset, collected from actual sales scenarios, showcases the practical application value of TGTSF.
Experimental results show that TGForecaster consistently demonstrates state-of-the-art performance across our datasets, affirming the soundness of the TGTSF task definition. Ablation studies reveal that without textual assistance, TGForecaster’s performance reverts to that of PatchTST (Nie et al., 2023), its time series encoder backbone, underscoring that the performance enhancement is driven by the additional information provided by textual data, not merely by a more sophisticated architecture.
Our contributions are summarized as follows:
  • We identify the roadblock of TSF as information insufficiency and propose TGTSF as a new forecasting approach that integrates textual data to enrich the models with external causal information and system knowledge.
  • We establish the first TGTSF benchmark containing four uniquely designed datasets.
  • We design a simple baseline model for TGTSF, TGForecaster. TGForecaster achieves state-of-the-art performance by utilizing textual information and effectively validates our proposed TGTSF task.
2 Insights and Motivation
2.1 Existing TSF Solutions Suffer from Information Insufficiency
Time series forecasting typically involves predicting future segments based on historical data, specifically forecasting subsequent time series segments from prior ones. However, the inherent sparsity of information within time series data makes achieving accurate forecasts challenging, or even impossible at times. Standard decomposition of time series data identifies three main components: trend, periodicity, and noise (Box et al., 2015). The noise component is intrinsically unpredictable. Trends, which are slowly changing patterns, are often impacted by external events that influence the underlying system, making it less predictable without external information. As a result, the periodicity component is usually the main source of reliable prediction.
Recent advancements in linear models have validated these challenges. For instance, the DLinear (Zeng et al., 2023) model utilizes a simple linear layer to effectively capture basic periodic patterns, outperforming most complex Transformer-based models. Similarly, FITS (Xu et al., 2023), designed to simply extract periodicity, achieves comparable or superior results to state-of-the-art (SOTA) methods using fewer than $10k$ parameters. Further research demonstrates that employing a closed-form forecasting matrix (Toner & Darlow, 2024), calculated from training data, can produce SOTA outcomes. These findings indicate that the minimal information in time series data allows models to recognize patterns with few or no learnable parameters, suggesting that training larger models with insufficient information can lead to severe overfitting.
The issue of information insufficiency extends beyond the data itself to the absence of external information. External factors like holidays, consumer sentiment, or climate changes can significantly alter patterns in time series data, affecting variables such as electricity usage, sales trends, or regional weather conditions. However, these factors are not contained within the time series data, hindering models from incorporating causal relationships. This deficiency can compromise model training, causing models to either learn overly simplified average shortcut or overfit on the training data, as shown in Fig. 1. For example, if a dataset frequently presents days without rain, the model may consistently predict no rain for future days. However, rainfall predictions are largely influenced by external factors like climate change, not included in the time series, leading models to opt for shortcuts that minimize overall loss by predicting the most common outcome.
Moreover, the lack of knowledge about the underlying systems exacerbates the problem of information insufficiency, especially in datasets with multiple time series channels. Without a deeper understanding of these systems or the characteristics of each channel, it is challenging to leverage such information for improved modeling performance. The indistinctiveness among channels complicates modeling efforts, making it difficult to accurately capture and model correlations.
These challenges underscore the need to integrate external information to address the fundamental limitations of time series forecasting methods.
2.2 RIN & Weight Sharing are Compromises under Information Insufficiency
Researchers have developed various methods to tackle information insufficiency in time series forecasting. One such method is Reversible Instance Normalization (RIN) (Kim et al., 2021), which normalizes each instance to align mean and variance across the dataset, as shown in Fig. 1. RIN effectively reduces distribution shifts caused by trends shifting or external events, forcing the model to learn a unified distribution. While it standardizes data inputs, RIN also strips away essential information such as trend intensity and amplitude variations, which diminishes the model’s capacity to detect relative biases or amplitude shifts beyond the training scope. However, RIN merely hides the issue of information insufficiency by neutralizing the effects of external events.
Figure 1: Left Panel: The traditional TSF model, when faced with distribution shifts due to external events, tends to learn an averaged shortcut for a smaller global loss, as shown in the leftmost figure. Conversely, the TGTSF task utilizes text inputs to eliminate the uncertainty and effectively generalize to these shifts, as depicted in the second figure. Right Panel: The rightmost figure illustrates the effect of text guidance on adapting to different channel distributions. Weight Sharing (WS) method helps the model capture only common knowledge among channels, while RIN addresses distribution shifts by normalizing instances, thus learning a larger portion of the distribution. TGTSF, however, leverages text to directly learn the entire distribution, enabling the model to generalize to a specific channel distribution with corresponding text as the condition.
Another strategy is weight sharing (Nie et al., 2023; Zeng et al., 2023; Xu et al., 2023). This technique treats all input channels as equivalent, leveraging common patterns across all channels to alleviate the problem of limited training samples. In other words, this approach enhances performance by learning from the shared periodicities across all the channels, typical within a single dataset. However, since the model is trained to minimize loss across all channel distributions with a single set of weights, it prioritizes common patterns, failing to capture unique distributions for each channel, as illustrated in Fig. 1, which may lead to averaged results. These average results further reveal that the model lacks sufficient information to distinguish between different channels effectively.
Despite these advancements, both methods provide only temporary solutions to the pervasive issue of information insufficiency. They do not resolve the core problem and still suffer from the limitations imposed by insufficient information.
2.3 Motivation for The Proposed TGTSF Framework
External factors significantly impact time series behavior such as trending and periodicity. A line of work apply and investigate how to embed the external information as a new channel, such as Autoformer (Wu et al., 2021), FedFormer (Zhou et al., 2022a) add auxilary channels to indicate the day-of-week or even the exact date. Temporal Fusion Transformer (TFT) (Lim et al., 2021) and TimeXer (Wang et al., 2024), however, focus on how to select the given variables and how to embed and fuse them with other channels. However, these methods only bring limited performance boost. Basically, they are still trying to squeeze out the information in the original time series dataset which contains limited information itself.
However, other forms of information such as news articles, social media updates, global climate change report are crucial for accurately predicting time series pattern in domains like sales and local weather report. These information are often in forms of text and hard to be quantified as numerical data alone. Moreover, the real-world application of forecasting frequently requires conditional predictions based on specific scenarios. For example, businesses may need to assess the impact of a marketing campaign on future sales, necessitating a deep understanding of external textual factors to forecast outcomes effectively.
To address these challenges, it is essential to include text as an additional input modality. This approach not only allows models to react to historical data but also to proactively adapt to potential future events and trends. It enhances forecasting by providing insights into the underlying dynamics and causal relationships often missing from time series data alone. Guiding time series forecasting with textual data also helps clarify inter-channel relationships, reflecting systemic correlations or domain-specific knowledge that are crucial for accurate predictions. This textual context enhances models’ ability to interpret complex data interactions, vital in environments where data channels are interconnected.
In summary, guiding time series forecasting with textual data addresses the critical issue of information insufficiency. This approach transcends the limitations of traditional TSF methodologies, enhancing model accuracy and effectiveness. By incorporating textual insights, forecasting becomes more robust and practically valuable for real-world applications.
3 TGTSF Task Formulation
We introduce a novel task within the domain of time series analysis, termed Text-Guided Time Series Forecasting (TGTSF), designed to address the prevalent issue of information insufficiency in time series forecasting.
As illustrated in Fig. 1, text guidance in TGTSF operates on two aspects. Firstly, channel descriptions serve as identifiers for each channel, aiding the model in distinguishing between them while learning shared features. These descriptions can also incorporate static knowledge about the underlying system, enhancing the model’s ability to recognize inter-channel correlations. Secondly, news messages provide dynamic and external insights into known (in training) or hypothesized (in inference) future events, which assist the model in adapting to event-driven distribution shifts.  [^1] In creating a TGTSF dataset, it is advisable to avoid directly generating the description out of the forecasting horizon time series pattern as news messages, as this could lead to information leakage. News messages should instead contain relevant, known information from other sources.
The TGTSF task is defined by three principal inputs: the time series data, news messages, and channel descriptions. The challenge lies in the fact that news items are not directly linked to specific time series channels. Consequently, the model must infer the impact of each news item on individual channels based on their descriptions. This setup not only capitalizes on dynamic events but also integrates domain knowledge via channel descriptions. Such integration is crucial for the model to comprehend and learn the spatial correlations among channels, thereby enhancing the forecasting accuracy with nuanced contextual understanding.
Formally, the TGTSF task seeks to model the potentially complex joint distribution of future values in a multi-channel sequence. This modeling involves the integration of historical time series data, textual information from news messages, and channel descriptions over a specified look-back window:
$P(X_{n:n+h}^{j}|(X_{n-L:n-1}^{j},News_{n:n+h},Des^{j})),j\in[1,c]$ (1)
where $L$ denotes the look-back window length, $h$ stands for the prediction length, the $j$ means the channel number, $c$ is the total number of channels.
4 TGForecaster: A Baseline Model for The TGTSF Task
To validate the TGTSF task, we developed TGForecaster, a streamlined transformer model designed for multimodal fusion with cross-attention. Illustrated in Fig. 2, this model harnesses textual information to enhance the accuracy and relevance of time series forecasting. It employs a reimplemented PatchTST encoder for processing time series data and leverages off-the-shelf, pretrained text models for pre-embedding textual inputs into vector sequences across the time dimension, allowing for effective modality fusion in the embedding space.
Figure 2: Architecture of the TGForecaster Model. TGForecaster integrates three primary inputs: time series data from a look-back window, news embeddings, and channel description embeddings. The texts are aligned to the forecasting horizon length. Text-side embeddings act as queries within the modality mixer layer to synchronize outputs with the forecast length. On the right panel, TGForecaster’s efficiency is highlighted by the text encoder’s cross-attention layer, which effectively associates news with specific channels. The token-wise decoder prevents overfitting in the final linear layer through inverse patching. Key parameters—number of patches ($PN$), maximum news items per batch ($M$), look-back window ($L$), and forecast horizon ($H$)—are streamlined for clarity.
A key innovation in TGForecaster is the integration of a cross-attention layer to achieve the ’text guided channel independent’. This layer is pivotal in fusing news content with channel descriptions. Taking cues from the attention mechanism’s original use in recommendation systems, we treat the news as both the key and value, with the channel descriptions serving as queries. Through this mechanism, we calculate the relevance of each news item to every channel. The resulting scores are then used to aggregate the news, culminating in a composite embedding for each channel. Another crucial aspect of our design is the modality fusion, achieved through a separate layer of cross-attention where text embeddings function as query, while temporal embeddings act as the key and value. This ensures that the output dimensions align naturally with the query which has the length of forecasting horizon. The attention maps for these two cross-attention layers are provided in Appendix E. These maps compellingly demonstrate that the layers perform in accordance with our design specifications.
As decoder, we use the token-wise decoding to avoid the overfitting on the last layer as observed in previous work (Lee et al., 2023). Each token is project back to its original patch length and weighted sumed to get the final forecasting result. Notably, TGForecaster does not incorporate the RIN. Our experiments, detailed in Appendix D, demonstrate that the inclusion of RIN degrades performance when sufficient information is already available.
This design strategy not only enables the integration of textual data into the time series forecasting model but also facilitates a direct comparison of the impact of text on forecasting accuracy.
5 TGTSF Benchmark Datasets
We have designed four TGTSF benchmark datasets across three categories: synthetic, captioned existing, and real-world datasets. We plan to release these TGTSF datasets along with captions, captioning scripts, and pre-embeddings, adhering to all applicable intellectual property constraints.
5.1 Synthetic Toy Dataset
The Synthetic Toy Dataset evaluates models’ ability to utilize textual information for time series forecasting. It features diverse patterns, including segments of sinusoidal waves with variations in frequency, amplitude, and trend. Textual descriptions resembling news reports precede each change point, outlining upcoming alterations to prompt the TGTSF model to adjust its forecasts accordingly. This setup tests the model’s effectiveness in mitigating distribution shifts and harnessing textual cues for improved forecasting accuracy.
5.2 Electricity-Captioned Dataset
The Electricity-Captioned Dataset builds on a commonly-used electricity utility dataset (Zhou et al., 2021) that tracks appliance usage in an office building, with daily patterns largely influenced by whether it is a workday. This dataset has been enhanced with straightforward, publically-known textual information such as the type of day (e.g., day of the week, public holiday, workday), using the channel name as the descriptor. This enhancement serves a dual purpose: to illustrate whether minimal textual data can improve prediction accuracy and to facilitate a direct performance comparison with traditional time series models.
5.3 Weather-Captioned Dataset
The Weather-Captioned Dataset is designed to overcome the limitations of commonly used datasets in time series forecasting (TSF), especially the lack of predictability and periodicity in variables like rainfall, wind speed, and direction. By incorporating external climate information, this dataset demonstrates how integrating textual data can address the challenges of information insufficiency in traditional TSF, enabling enhanced control over forecasting outcomes.
Originally limited to a single year (Zhou et al., 2021), we have expanded the weather dataset to encompass a decade of detailed weather data from 2014 to 2023, sourced consistently from the same weather station. The Weather-Captioned dataset now includes weather forecasting reports from a publicly accessible service, detailing the climate conditions in Jena, Germany, where the weather station is approximately located. These reports, updated every six hours and daily, cover a range of parameters including weather conditions, temperature, humidity, wind speed, and direction, providing a comprehensive overview for our analysis. We directly use the channel name as channel description.
Table 1: Example caption of the Weather-Captioned dataset.
Topic Example
Month & Time of the Day It’s the early morning of a day in January.
Overall Weather The current weather is clear.
Weather Trend in next 6h The weather is expected to remain clear.
Temprature Trend in next 6h The temperature is showing a mild drop.
Wind Speed & Direction There is Light Breeze from NNW.
Atmosphere Pressure Level The atmospheric shows Average Pressure.
Humidity Level The air is very humid.
To transform this quantitative data into actionable insights for TSF, we employed GPT-4 to summarize each forecasting report into a set of seven thematic sentences, as demonstrated in Tab. 1. GPT-4 is prompted to avoid specific numerical details from the original reports to prevent information leakage and to simplify the model’s learning process. To further enrich the dataset, GPT-4 generated three distinct summary versions for each report, resulting in $21873^{7}=$ 2187 possible unique textual captions for each time step. The full details of the prompt used for generating these summaries are provided in the appendix.
We divided this extensive dataset into two subsets for detailed analysis: Weather-Captioned-Medium, covering data from 2014 to 2018, and Weather-Captioned-Large, which includes the entire dataset spanning ten years. This dataset encompasses over 525,600 weather data records across 21 different channels, with more than 85,000 unique sentences. This vast corpus provides millions of potential training samples for captioned time series forecasting. Further information is detailed in the appendix.
5.4 Steam Game Dataset
A key application of TGTSF is in sales forecasting and decision-making. To illustrate this, we compiled data on online gamer counts from several of the most popular games on Steam, the leading online video game distribution platform. This dataset also includes records of all game updates and events. The time series exhibits a weekly periodicity, with gamer activity typically peaking on weekends. Announcements from game developers frequently trigger spikes in activity, reflecting player enthusiasm for new content. However, the varied management strategies of game developers and the diverse reactions of gamers to updates introduce significant distribution shifts, even when textual information is considered. This variability makes the Steam Dataset the most challenging and intricate dataset within the TGTSF framework.  [^2] Due to intellectual property constraints, we may not be able to directly release this dataset. Detailed instructions are available in the appendix for researchers interested in replicating or extending our data.
6 Experiments
We evaluate the TGForecaster model across four datasets to demonstrate the feasibility of the TGTSF task on the proposed datasets. As discussed in Section 5, each dataset is uniquely designed to test the task from different perspectives.
Baseline Models
TGForecaster is benchmarked against state-of-the-art (SOTA) methods including DLinear, FITS, PatchTST, iTransformer, and Time-LLM (Zeng et al., 2023; Xu et al., 2023; Nie et al., 2023; Liu et al., 2023; Jin et al., 2023). We specifically contrast it with the linear-based models DLinear and FITS to illustrate how TGForecaster surpasses periodicity-focused models. Comparisons with PatchTST highlight how textual information can enhance forecasting performance, even when using the same time series encoder. Time-LLM "reprograms" the LLaMa2 (Touvron et al., 2023), a large language model, which gives it capability of understanding text information. We use it as a baseline of multi-modal model. Specific settings differences, if any, are noted accordingly.
6.1 Evaluation on Toy Dataset
Experiment Settings
All of the models are following the same experimental setup with prediction length $H\in\{14,28,60,120\}$ and LBW length $T=60$.
Statistical Results
Table 2 presents the performance comparison of various models on the toy dataset, with TGForecaster significantly outperforming all baseline models. Specifically, TGForecaster achieved an 80% improvement in Mean Squared Error (MSE) over the best performing transformer-based model, and a 96% increase over DLinear. These results underscore the considerable benefits of integrating textual guidance in TSF. Notably, a version of TGForecaster without news data, which lacks the ability to utilize auxiliary textual information, demonstrated substantially lower performance. This underscores the essential role of text in enhancing forecast accuracy.
Table 2: Forecasting result on toy and Electricity-Captioned dataset in MSE. The best result is highlighted in bold and the second best is highlighted in underline.
Dataset Pred. Len. TGForecaster
TGForecaster
w/o News
FITS DLinear PatchTST iTransformer TimeLLM
14 0.003 0.006 0.282 0.151 0.006 0.136 0.231
28 0.008 0.018 0.692 0.297 0.029 0.295 0.382
60 0.020 0.052 0.909 0.442 0.075 0.494 0.551
Toy 120 0.027 0.102 0.883 0.632 0.168 0.747 0.788
Pred. Len. TGForecaster
TGForecaster
lbw 120
FITS DLinear PatchTST iTransformer TimeLLM
96 0.124 0.127 0.134 0.140 0.130 0.148 0.131
192 0.144 0.146 0.149 0.153 0.149 0.162 0.152
336 0.160 0.164 0.165 0.169 0.166 0.178 0.160
Elec-c /Elec 720 0.193 0.200 0.203 0.204 0.210 0.225 0.192
Case Study
Figure 3 illustrates a segment of the toy dataset where the frequency changes within the forecasting horizon. In this visualization, the PatchTST model maintains the frequency observed in the look-back window, indicating an inability to adapt to new frequencies. Similarly, DLinear displays a collapsed pattern. Notably, without news input, TGForecaster shows the same behavior of PatchTST, suggesting that it relies solely on its time series encoder in the absence of textual cues. The models without external data also highlights the decreasing amplitude of predictions, a tendency to revert to safer, average predictions when faced with uncertainty in the far future—an issue known as the ’average shortcut’ phenomenon discussed in Section 2.1. Conversely, with the integration of external text information eliminating uncertainty, TGForecaster adeptly adapts to new frequencies at the appropriate moments, demonstrating the substantial benefits of incorporating textual data into the forecasting process.
Figure 3: Visualization on the toy dataset. Look-back window is 60. Orange line for prediction and blue line for groundtruth. This segment shows two frequency change points in the forecasting horizon. With the help of external text information, TGForecaster can accurately adapt to new frequency at the correct position, however other conventional time series models fail to do so.
6.2 Evaluation on Electricity-Captioned Dataset
Experiment Settings
We follow the experiment settings in previous works as follows: forecasting horizon $H\in\{96,192,336,720\}$, look back window length of 288. For fair comparison, we directly compare with the results report in the baseline original paper. And the TGForecaster is trained with the captioned version. On this dataset, we also test the impact of a shorter look-back window on the TGForecaster. We will report accordingly in Tab. 2.
Statistical Results
As depicted in Tab. 2, TGForecaster demonstrates SOTA performance on the Electricity dataset, particularly effective at shorter forecasting horizons using minimal textual cues. However, TimeLLM slightly outperforms TGForecaster over longer forecasting periods. Given TimeLLM’s backbone is a large language model that also incorporates textual information, its edge may stem from its use of basic date and statistical data within its text inputs.
6.3 Evaluation on Weather-Captioned Dataset
Experiment Settings
We follow the experiment settings in previous works as follows: forecasting horizon $H\in\{96,192,336,720\}$, look back window length of 360. We trained all other models on the Weather-Medium and -large dataset with their setting on the original weather dataset accordingly. And the TGForecaster is trained with the captioned version. We will report accordingly in Tab. 3.
Table 3: Forecasting result on Weather-Midium and -Large dataset in MSE. The best result is highlighted in bold and the second best is highlighted in underscore.
Dataset Pred. Len. TGForecaster FITS DLinear PatchTST iTransformer TimeLLM
96 0.182 0.248 0.294 0.252 0.267 0.294
192 0.205 0.297 0.340 0.304 0.327 0.342
336 0.235 0.354 0.393 0.364 0.404 0.393
Weather -Medium 720 0.281 0.430 0.456 0.439 0.495 0.461
96 0.410 0.436 0.487 0.464 0.456 -
192 0.438 0.524 0.568 0.567 0.578 -
336 0.455 0.601 0.644 0.644 0.698 -
Weather -Large 720 0.497 0.692 0.725 0.745 0.832 -
Statistical Results
As demonstrated in Tab. 3, TGForecaster significantly outperforms other models across both the Weather-Medium and Weather-Large datasets, substantiating the efficacy of incorporating external text information in addressing the information insufficiency inherent in TSF models. Notably, despite its simplicity, the FITS model performs exceptionally well, second only to TGForecaster. This suggests that simply adding more training time series to complex models does not necessarily lead to improved performance, as more time series does not necessarily contain more information. The results highlight that strategic integration of textual data can provide a more substantial performance boost than merely increasing the quantity of time series data. Further, we also report the channel-wise performance in Appendix B. We notice a groundbreaking performance boost of over 60% on certain channels which were not predictable with historical time series data alone.
Table 4: Ablation result on Weather-Captioned-Medium in MSE.
Pred.
Len.
Openai
512
MiniLM mpnet
MiniLM
w/o Des.
MiniLM
w/o News
96 0.182 0.186 0.196 0.209 0.249
192 0.205 0.214 0.216 0.260 0.302
336 0.235 0.232 0.251 0.302 0.359
720 0.281 0.272 0.291 0.356 0.432
Ablation Study
We conducted ablation studies to evaluate the impact of different embedding models and the integration of textual inputs on the performance of TGForecaster. We tested three embedding models: OpenAI Embedding (ope, ), paraphrase-MiniLM-L6 (Wang et al., 2020), and all-mpnet-base (Song et al., 2020) [^3] Sentence Transformer: https://huggingface.co/sentence-transformers. The results, presented in Tab. 4, indicate minimal performance differences between the embedding models. Notably, the removal of channel descriptions led to a significant performance decrease, underscoring the model’s reliance on this feature for distinguishing between channels. Similarly, omitting news text resulted in performance dropping to levels comparable to the baseline PatchTST model, confirming that the improvements in forecasting accuracy are primarily driven by the inclusion of external textual information. This observation validates our hypothesis that textual data plays a crucial role in compensating for information deficiencies in traditional time series forecasting.
Visualization and Controllability Test
Fig. 4 visualizes three channels from the Weather-Caption-Medium dataset, full result see Appendix C. The first channel, atmospheric pressure (p), which is influenced by regional climate conditions, typically exhibits slowly changing trends that are challenging for TSF models to predict due to their subtle fluctuations without periodicity. However, with the integration of external information, TGForecaster accurately predicts these trends, whereas PatchTST tends to predict a constant average value, failing to capture the gradual changes.
Figure 4: Visualization three channels on the 15000th test sample of Weather-Caption-Medium dataset. Blue line for ground truth, Red line for TGForecaster, Green line for PatchTST and Orange line for TGForecaster with swapping the news on the second and forth forecasting day.
The second channel, time of raining, measured in seconds per 10 minutes, presents unique challenges: it lacks periodicity, displays no causality, and often exists in a binary state. Traditional models like PatchTST often fail, typically predicting no rainfall - the averaged shortcut. TGForecaster adjusts its forecasts based on text prompts about upcoming rain. However, it sometimes misses predictions when the text weather forecasting report does not perfectly align with actual weather events, as noted with the missed second rainfall period.
The third channel, SWDR (solar radiation), directly reflects the amount of solar power reaching the ground. TGForecaster accurately forecasts SWDR, even though these are not directly mentioned in the news texts. It effectively infers inter-channel dependencies, enabling it to predict shifts in SWDR associated with weather changes, in contrast to PatchTST’s basic waveform predictions.
Additionally, a controllability test involving the swapping of news inputs for the second and fourth days illustrates TGForecaster’s responsive accuracy: it predicts rain on the fourth day and clear conditions on the second, aligning forecasts with the altered news data. This test underscores the model’s adaptability and the effectiveness of text-guided forecasting.
6.4 Evaluation on Steam-100 Dataset
We evaluate the performance of the TGForecaster on our Steam-100 dataset, utilizing an input window of 60 days and an output horizon of 14 days. The findings indicate that the TGForecaster outperforms baseline models such as PatchTST, achieving a performance enhancement of over $12.6\%$. This superior performance is consistently observed across over $59.6\%$ of all games, ranking as the best among all evaluated methods. Given the significant stylistic variations among different game developers and the potential risk of temporal distribution shifts, traditional time series forecasting methods often struggle to capture the commonalities in temporal features. In contrast, the TGForecaster leverages textual information to significantly augment its predictive capabilities. We report the full result in the Appendix G.
7 Conclusions and Discussion
This paper addresses a critical roadblock in time series forecasting: information insufficiency. We introduced Text-Guided Time Series Forecasting (TGTSF), a new approach that integrates textual cues to enrich the models with external information and system knowledge. We developed and released four TGTSF datasets, each crafted to validate different aspects of the task and model. Our straightforward yet effective TGForecaster model demonstrates that textual guidance can significantly enhance time series modeling by mitigating the average predictions typically resulting from information scarcity.
While the TGForecaster effectively validates the TGTSF task, it does not fully comprehend the semantics of the text, such as extracting correlations among channels automatically. Future work will focus on advancing the model’s semantic understanding and its ability to autonomously discern intricate relationships within the data.
Appendix A Data and Code Availability
Given the extensive size of the TGTSF dataset, especially the ones with pre-embeddings, an immediate anonymous release may not be feasible. Therefore, we plan to make the full datasets publicly available ASAP.
Additionally, the script used to create the toy dataset will be released alongside our code base. You can access the code repository at: https://github.com/VEWOXIC/TGTSF.
We will update the github repo here later and will update the dataset in it later.
Appendix B Channel-wise performance on Weather-Captioned Dataset
The difficulty in predicting each channel varies, therefore, we present channel-wise performance in Table 5. The results demonstrate that TGForecaster, with the aid of external textual climate reports, significantly enhances forecasting accuracy across all channels. Notably, the model achieves over a 60% performance improvement in channels such as atmospheric pressure (p (mbar)), relative humidity (rh(%)), and vapor pressure deficit (VPdef (mbar)), which typically cannot be predicted reliably using historical time series data alone. The integration of external text cues has led to groundbreaking improvements in forecasting these parameters.
However, the wind velocity channel shows minimal variance in performance across all models, each achieving similar results with slight losses. This phenomenon is attributed to the presence of extreme values in this channel, which, after normalization, diminish the impact of more typical values on the overall gradient. Consequently, all models struggle to learn detailed patterns in this channel due to the reduced contribution to the global gradient.
Another noteworthy observation is that while TGForecaster is capable of predicting rainfall—unlike models that default to predicting near-zero averages—the performance improvement in these channels is modest. This is because rainfall is relatively scarce in this dataset, leading to large losses when rain is inaccurately predicted at the wrong times. Conversely, predicting the average value results in a smaller overall loss. This tendency explains why other models often opt for the average, avoiding the complex task of learning rainfall patterns. Nevertheless, accurate rainfall forecasting remains crucial in meteorological applications, underscoring our commitment to enhancing predictive accuracy in this area. The same principle also applies to other channels.
Table 5: Channel wise performance on Weather-medium dataset in MSE. The best is highlighted in bold and the second best is highlighted in underline.
Channel TGForecaster FITS DLinear PatchTST iTransformer IMP.
p (mbar) 0.1365 0.8637 0.8238 0.9301 1.0320 83.43%
T (degC) 0.1889 0.2924 0.3329 0.2964 0.3233 35.40%
Tpot (K) 0.1829 0.3163 0.3525 0.3225 0.3533 42.18%
Tdew (degC) 0.3467 0.4043 0.4085 0.4082 0.4258 14.25%
rh (%) 0.2479 0.6541 0.6788 0.6997 0.8185 62.10%
VPmax (mbar) 0.2369 0.3500 0.3984 0.3521 0.4086 32.31%
VPact (mbar) 0.2998 0.3404 0.3534 0.3515 0.3845 11.93%
VPdef (mbar) 0.2835 0.6384 0.6968 0.6744 0.8038 55.59%
sh (g/kg) 0.2995 0.3434 0.3562 0.3557 0.3896 12.78%
H2OC (mmol/mol) 0.2996 0.3434 0.3562 0.3556 0.3894 12.75%
rho (g/m³) 0.1926 0.3909 0.4119 0.4182 0.4535 50.73%
wv (m/s) 0.0002 0.0002 0.0002 0.0002 0.0003 0.00%
max. wv (m/s) 0.0004 0.0005 0.0004 0.0005 0.0006 0.00%
wd (deg) 0.7270 1.1605 1.1295 1.1344 1.2735 35.64%
rain (mm) 0.6824 0.6905 0.7167 0.6891 0.6998 0.97%
raining (s) 0.7900 0.8735 0.9379 0.8591 0.9942 8.04%
SWDR (W/m²) 0.1828 0.3084 0.3856 0.2967 0.3776 38.39%
PAR (umol/m²/s) 0.1773 0.2840 0.3588 0.2704 0.3473 34.43%
max. PAR (umol/m²/s) 0.1975 0.2599 0.3195 0.2632 0.3226 24.01%
Tlog (degC) 0.1774 0.2802 0.3260 0.2806 0.3290 36.69%
CO2 (ppm) 0.2600 0.2716 0.2812 0.2618 0.2760 0.69%
Avg. Loss 0.2814 0.4317 0.4583 0.4391 0.4954 34.82%
Appendix C Visualization Weather-Captioned Dataset
We provide the full visualization as Fig. 5. The TGForecaster shows great performance across all the channels. Even very hard ones such as Wind dir. It can also model the time series that totally independent with the weather such as the CO2 channel.
Figure 5: Full visualization all channels on the 15000th test sample of Weather-Caption-Medium dataset. Blue line for ground truth, Red line for TGForecaster, Green line for PatchTST and Orange line for TGForecaster with swapping the news on the second and forth forecasting day.
Appendix D Weather Results w. w/o. RIN
We compared the performance of models with and without Reversible Instance Normalization (RIN) on the weather-captioned-medium dataset, focusing on a 720-hour forecasting horizon. The model with RIN enabled achieved an MSE of 0.3428, whereas the model without RIN achieved a lower MSE of 0.2814. Results visualized in Fig. 6 show that the RIN-enabled model exhibits significant biases in many channels, particularly those with gradual trend shifts. This occurs because RIN removes the bias term from all instances, leaving the model unable to recognize relative bias and trend values. For instance, with RIN, temperature patterns in winter and summer are treated similarly, ignoring the typically higher and more variable temperatures in summer. Additionally, we noted pronounced shifting behavior coinciding with changes in captions, suggesting that the absence of bias information leads the model to over-rely on textual prompts, compensating for the missing data.
Figure 6: Full visualization all channels on the 15000th test sample of Weather-Caption-Medium dataset. Blue line for ground truth, Red line for TGForecaster without RIN, Green line for TGForecaster with RIN enabled.
Appendix E Attention Map Visualization on Weather-Captioned
We further visualize two cross-attention blocks to further investigate the TGForecaster. You are strongly advised to check the Tab. 1, Appendix K.3 and Fig. 5 while reading this part.
Figure 7 illustrates the attention map of the "text-guided channel independent" cross-attention block in the text encoder across three layers. In the first layer, attention is predominantly focused on the first sentence, which specifies the month and time. This sentence is crucial as it provides temporal context that significantly impacts the prediction of both daily and annual periodicity. While other sentences receive moderate attention, the sixth sentence, which describes atmospheric pressure as detailed in Table 1, consistently receives no attention across all channels.
In the second layer, however, there is a notable shift in attention dynamics. All channels, particularly channel 0, show intense focus on the sixth sentence. According to the channel definitions in Appendix K.3, channel 0 directly corresponds to atmospheric pressure. Channels 10 and 20, which are related to air density and CO2 concentration respectively—factors closely associated with pressure—also display relatively high attention scores. This suggests that the TGForecaster is capable of discerning the underlying relationships among the channels.
The separation of attention focus between the first and second layers suggests that the influence of atmospheric pressure on the model’s predictions is independent of time. In the third layer, a diversity of attention patterns emerges; channel 0 focuses exclusively on the sixth sentence, while other channels predominantly attend to the first sentence.
Since we take the output of previous layer as query and input news embeddings as key and value, the information lies in the news are progressively added to the channel embeddings. Thus, the model can focus on different perspective in separate cross attention layers.
(a) Layer1
(b) Layer2
(c) Layer3
Figure 7: Attention map of the "text guided channel independent" cross attention block on weather-captioned dataset, on on the 15000th test sample of Weather-Caption-Medium dataset. We use three cross attention block. The vertical axis stand for channels and horizon stand for the 7 sentences of the weather report summary.
Figure 8 presents the attention map of the modality mixer layer cross attention block in the weather-captioned dataset. The map, averaged across three cross attention layers, illustrates distinct patterns of attention for each channel. This diversity underscores the TGForecaster’s ability to adaptively extract time series embeddings tailored to the unique distribution characteristics of each channel, facilitated by textual inputs.
Notably, the channels for SWDR, PAR, and max.PAR display clear periodic patterns in their attention maps, aligning with observations from waveform visualizations. These patterns suggest that the TGForecaster effectively captures and utilizes periodic information from these environmental variables.
Furthermore, the channels labeled rain and raining show a particularly interesting behavior; they assign significantly higher attention scores to the exact time periods of rainfall within the look-back window. This behavior indicates that the TGForecaster is adept at identifying and prioritizing crucial temporal events specific to each channel, further enhancing its forecasting accuracy by focusing on relevant patterns where needed. This level of detail in attention allocation demonstrates the model’s capability to integrate contextual cues from textual data and further guide the time series forecasting.
Figure 8: Attention map of the modality mixer layer cross attention block on weather-captioned dataset, onon the 15000th test sample of Weather-Caption-Medium dataset. The attention map is averaged across three cross attention layers. We plot the attention map for each channel. The vertical axis stand for output time series patches and the horizon stand for the input time series patches embedding from PatchTST backbone.
Appendix F Comparison with More Baselines on Electricity-Captioned
We further compare with more baselines on Electricity-Captioned, including Autoformer, Fedformer, Informer, FiLM and TimesNet (Wu et al., 2021; Zhou et al., 2022a, 2021, b; Wu et al., 2023). TGForecaster shows dominant superior performance across these baselines, as shown in Tab. 6.
Table 6: The comparison on Electricity dataset with other baselines. Best is marked in bold and the second best is marked in underline.
Pred. Len. TGForecaster TGForecaster_120 Autoformer Fedformer Informer FiLM TimesNet
96 0.124 0.127 0.201 0.188 0.274 0.154 0.168
192 0.144 0.146 0.222 0.197 0.296 0.164 0.184
336 0.16 0.164 0.231 0.212 0.3 0.188 0.198
Elec 720 0.193 0.209 0.254 0.244 0.373 0.236 0.22
Appendix G Full Result on Steam-100 Dataset
We show the comparison on Steam-100 Dataset in Tab. 7, and Tab, 8 with PatchTST. We use de-normed MAE as metric since the base volume of players of each game varies drastically, using normed metrics can lead to unfair comparison. The pretrain indicate that the model is jointly trained on all the games and each game is labeled by the channel discription. The Gnorm indicate that we apply the global normalization to preserve the player variation mentioned before. But it seems bring limited boost.
Table 7: Full result on Steam-100 dataset in de-normalized MAE. The best result is shown in green shaded bold font. The ones with performance boost over 10% is marked in red.
game_id TGTSF_pretrain
TGTSF_pretrain
_Gnorm
TGTSF PatchTST IMP/%
10 781.2257 724.5196 849.8968506 804.757019 0.099704
240 372.887 374.03445 400.0899353 421.2349243 0.114777
440 10216.276 10816.979 10694.47852 10828.37891 0.056528
550 4110.673 4437.846 4336.002441 4587.186035 0.103879
570 30179.111 30914.955 31916.02344 32031.38477 0.057827
620 674.1199 789.6992 799.2894897 710.8320313 0.051647
730 51687.87 50937.348 52638.79297 51069.08984 0.00258
3590 687.4657 775.8215 734.3128662 743.4817505 0.075343
39210 3310.422 3719.8003 3388.064453 3420.974609 0.032316
105600 4415.245 4899.586 4464.801758 4513.034668 0.021668
107410 1597.8743 1548.8544 1411.974243 1355.757813 0
214950 331.97876 350.3743 328.4424744 324.0109863 0
218620 5576.539 5714.511 5650.027832 5818.570313 0.041596
221100 2574.3164 2713.0063 3260.266113 2742.404053 0.061292
222880 50.29491 138.43388 61.51412582 59.99531174 0.161686
227300 3224.443 3356.0312 3418.429199 3388.108398 0.048306
230410 5307.4634 5733.2676 5856.409668 6040.648926 0.121375
231430 441.79767 377.2723 581.7993774 713.0888062 0.470932
232050 5.8684874 140.72949 6.452753067 6.638870239 0.116041
236390 4528.812 4847.2886 4787.857422 4680.506348 0.03241
236850 1197.9114 1308.7864 1169.570313 1191.723145 0.018589
242760 5888.873 6968.5566 6002.225586 6160.327148 0.044065
244210 657.4606 672.32324 676.5147095 658.0761108 0.000935
250900 895.2101 886.4052 1012.618652 892.572937 0.00691
251570 4304.9175 4886.079 4088.169922 4352.344727 0.060697
252950 1971.4385 1928.533 1971.546509 2054.135742 0.061146
255710 2154.8572 2082.0603 2231.175781 2078.98584 0
270880 789.74634 748.85596 746.2268677 760.31073 0.018524
271590 9292.364 9546.938 10758.82422 9438.995117 0.015535
275850 2671.1484 3121.9731 3179.639404 3118.741699 0.143517
281990 2094.3948 2404.5767 3269.309326 2315.452881 0.095471
284160 929.92413 903.6123 956.4987183 908.8114014 0.005721
289070 4861.033 5243.972 4706.715332 4633.394531 0
291550 1339.3384 1323.4893 1290.662231 1324.343018 0.025432
292030 3181.2307 3723.4417 3607.125732 3500.928223 0.091318
294100 2123.2292 2150.6 1990.630981 2074.705322 0.040524
304930 5698.3926 5597.64 5430.058105 5824.242188 0.06768
306130 1727.7567 1927.9446 1787.971802 1765.812744 0.021552
322170 987.1338 878.4059 902.7176514 910.0273438 0.034748
322330 4585.129 5155.708 4900.762207 4916.681152 0.067434
346110 6389.454 7178.1084 7004.43457 6871.126953 0.070101
359550 5251.984 5247.6694 5184.080566 5156.916016 0
364360 78.73614 182.35359 89.08701324 88.43521118 0.109674
365590 158.19469 229.93842 173.8761902 180.6734314 0.124416
374320 557.42993 630.0981 659.1234131 655.4638672 0.149564
377160 1733.9108 1486.6573 1814.430298 1725.577515 0.138458
381210 5498.251 5568.4336 5691.748047 5964.625488 0.07819
386360 1135.4182 1231.7946 1126.588989 1125.822021 0
394360 2788.8633 2674.6377 2861.023682 2725.428223 0.018636
413150 3061.9893 3204.8018 2831.16748 2713.414551 0
427520 923.96985 772.3062 803.0150146 775.4301758 0.004029
457140 1159.0262 1217.2543 1133.615601 1107.060547 0
489830 2071.2698 2067.1377 1855.288574 1809.03479 0
493520 270.12488 337.2064 311.7146606 314.7081604 0.141665
513710 1662.712 1732.691 2458.639648 2096.899658 0.207062
526870 1714.113 1704.2557 2002.178223 2031.111084 0.160924
Table 8: Cont. Full result on Steam-100 dataset in de-normalized MAE. The best result is shown in green shaded bold font. The ones with performance boost over 10% is marked in red.
game_id TGTSF_pretrain
TGTSF_pretrain
_Gnorm
TGTSF PatchTST IMP/%
529340 2037.8783 3383.1628 4038.390381 3701.234863 0.449406
548430 3215.861 3539.4717 3559.275391 3622.380615 0.112224
552500 2653.9133 2886.299 3560.82251 4062.790527 0.346776
552990 5025.465 5687.0366 3085.568848 5004.588379 0.383452
578080 21942.736 20576.87 24925.38086 21725.19727 0.052857
582010 2820.5447 3062.556 3161.88623 3088.219727 0.086676
582660 1543.2185 1686.1365 1597.595703 1613.270874 0.043423
646570 1219.2263 1304.4564 1372.947632 1420.136963 0.141473
648800 1754.793 2505.1694 2167.084717 2164.018311 0.189104
739630 3801.0945 4393.0825 4291.587891 4325.070313 0.121149
761890 1032.8406 1291.7933 1046.670288 1021.805847 0
814380 1362.8702 1614.8129 1658.933594 1566.417725 0.129945
892970 3217.5664 2532.8167 3313.418701 3676.412598 0.311063
960090 1899.8359 2076.6377 2291.049316 2292.431396 0.171257
1085660 16838.436 18822.541 18422.99219 18631.20313 0.096224
1091500 13698.44 14906.672 15611.44824 16007.56543 0.144252
1172470 33071.402 39898.113 37495.97266 37135.38672 0.109437
1172620 2595.5396 3012.9434 2990.058838 3043.287109 0.147126
1222670 3100.3125 2550.8154 3211.132813 3179.268799 0.197672
1238810 1900.4874 1945.948 2186.150391 2113.537598 0.100803
1238840 1193.3369 1223.7898 1373.025757 1397.660156 0.14619
1293830 1417.9685 1575.0452 1349.839478 1393.606567 0.031406
1326470 1735.4362 1926.2095 2924.927979 2730.827881 0.364502
1361210 4610.036 4732.349 9080.219727 6853.158203 0.327312
1454400 809.03754 769.7717 941.6995239 1514.508057 0.491735
1623660 576.48615 741.7087 850.6308594 978.7790527 0.411015
1665460 1282.9064 1174.2739 1324.958374 1345.970093 0.127563
1677740 1610.0262 1459.8435 1413.402588 1450.460693 0.025549
1811260 4966.4424 9192.259 8108.043457 7732.84668 0.357747
1868140 1495.137 1424.7719 12159.14551 9878.760742 0.855774
1919590 1274.8295 2378.603 6667.467773 1967.391602 0.35202
1938090 10918.149 10952.329 14469.45508 14213.95605 0.231871
1948980 534.38995 725.96063 1041.738037 1054.809937 0.493378
Best_count 53 17 9 10 0.126324
Appendix H Error Bar & Critical Difference Diagram
We run the experiments on Toy and Electricity for five times with different randomly chosen random seeds. And Weather-Medium for three times because of the large amount of data can result in very long training time on our devices. We report the mean and standard deviation as follows with comparison with FITS, the most stable model.
As Tab. 9 indicate, TGForecaster shows stable performance across the benchmark. Even with extreme condition, it still maintains superior performance. It worth note that, we thought the relative large variance on weather dataset is caused by the different combination of the text description. But the FITS also shows large variance on this dataset which indicate it is hard to converge on this dataset.
We generate the critical difference plot on our result of four datasets (toy, Electricity, Weather-Medium, Weather-Large) with the default alpha as 0.05 as shown in Fig. 9. TGForecaster’s placement at the top of the critical difference plot, without intersecting with other lines, demonstrates its consistent and superior performance in terms of MSE compared to the other models. It indicates that with the help of external textual information, TGForecaster can handle complicated datasets.
Figure 9: The Critical Difference Plot on the TGForecaster and other baselines with alpha=0.05.
Table 9: The mean and std of TGForecaster on the three dataset in metrics of MSE.
Datasets TGForecaster FITS
Toy 0.027±0.001 0.883±0.000
Electricity 0.193±0.004 0.203±0.001
Weather-Medium 0.281±0.008 0.430±0.011
Appendix I Toy Dataset Details
I.1 Data source
We directly generate this dataset with sinusoidal wave that randomly changes frequency. Before each changing point, we add 10 captions as ’Channel 1 will change to frequency x in y timesteps.’ After each changing point, we add 5 captions as ’Channel 1 will keep steady with frequency of x.’ In other timesteps, we caption it as ’The waveform will go steady.’
We will publish this dataset with CC BY-NC-SA 4.0 licence.
I.2 Dataset Structure
Toy/
|-- caption_emb/   // Pre-embeddings
|   |-- timestamp1.npy
|   |-- timestamp2.npy
|-- channel_1.npy     // channel description embedding data
|-- toy.csv     // time series data
Appendix J Electricity-Caption Details
J.1 Data source
We caption the day of week with the given time stamp. But we somehow find the original time stamp is incorrect. Instead of the year of 2016, it should be collected in year 2012. Without knowing the exact location of this building, we cannot identify the specific public holiday. We then uses channel 319, which shows obvious patterns of workday and holiday as indicator, when the average value lower than a specific value, we caption it with public holiday.
We will publish this dataset with CC BY-NC-SA 4.0 licence.
J.2 Dataset Structure
Elec-Captioned/
|-- caption_emb/   // Pre-embeddings
|   |-- Monday.npy
|   |-- Tuesday.npy
|   |-- Channel_1.npy
|   |-- Channel_2.npy
|-- electricity.csv     // time series data
Appendix K Weather-Caption Details
K.1 Data source
We get the weather time series data from: https://www.bgc-jena.mpg.de/wetter/ and weather report from https://www.timeanddate.com/weather/germany/jena/historic. We will publish this dataset with CC BY-NC-SA 4.0 licence since the data source forbids commercial use.
K.2 Prompt
We generate the summary of climate report using following prompt on GPT4.
You are a professional weather forecast message writer. You are provided with the weather
forecasting results in the next 6 hours and you should transcribe it as readable text.The
weather forecasting results are given in json string format. One is the coarse grained
weather of the next 6 hours and the fine grained json string contains the weather forecast
every half hour in these 6 hours.
You are suppose to summarise the weather forecast message in the following aspects, each
aspect should be a sentence or a phrase:
1. Time of Day, Month, e.g. "It’s the early morning of a day in December."
2. Current overall Weather Condition, you may use the term in coarse grained information,
e.g. "The current weather is clear."
3. Weather Trend in the 6 hours, you may summarise this according to the fine grained
information, e.g. "The weather is expected to remain clear." / "Rain is expected soon."
4. Temperature Trend in the 6 hours, you may summarise this according to the fine grained
information, e.g. "The temperature is showing a mild drop."
5. Wind Speed and Direction, you may summarise this according to the fine grained
information, e.g. "There is Light Breeze from NNW."
6. Atmospheric Pressure, describe the pressure of the atmosphere, e.g. "The atmospheric
shows very Low Pressure."
7. Humidity, describe the humidity of the atmosphere, e.g. "The humidity is very high."
The summary do not have to be very detailed, but should be clear and concise.
Note that, during summarization, you should NOT include the exact values of the weather
forecast, but only the trends and conditions. You should follow the following ranking
instructions:
1. Time of Day: $0$ 06:00 -> Early Morning, 06:00 - 12:00 -> Morning, 12:00 - 18:00
-> Afternoon, 18:00 - 24:00 -> Evening
2. Wind Direction: you should convert the wind direction of degrees to N, E, S, W, NE, SE,
SW, NW, NNE, ENE, SSE, WSW, NNW, ESE, SSW, WNW.
3. Wind Speed: you should convert the wind speed to: Less than 20 km/h -> Light Breeze,
20 to 29 km/h -> Gentle Breeze, 30 to 39 km/h -> Moderate Breeze, 40 to 50 km/h -> Fresh
Breeze, 51 to 62 km/h -> Strong Breeze, 63 to 74 km/h -> High Wind, 75 to 88 km/h -> Gale,
89 to 102 km/h -> Strong Gale, Over 102 km/h -> Storm.
4. Atmospheric Pressure: you should convert the atmospheric pressure to: (<990 mbar)
-> Very Low Pressure, (990-1009 mbar) -> Low Pressure, (1010-1016 mbar) -> Average
Pressure, (1017-1030 mbar) -> High Pressure, (>1030 mbar) -> Very High Pressure.
5. Humidity: you should convert the humidity to: (<30%) -> Very Dry, (30-50%) -> Dry,
(51-70%) -> Average Humidity, (71-90%) -> Humid, (>90%) -> Very High Humid. You
may change this to more oral expression.
6. Trend: you may use "increase", "decrease", "remain", "steady", "Go up/down"...
to describe the trend of the weather condition, temperature, wind speed,
atmospheric pressure, and humidity. You may change the expression.
Note that, the unit of the weather forecast may not provided, you should use
the following units:
 Temperature: Celsius, Wind Speed: km/h, Atmospheric Pressure/barometer: mbar,
 Humidity: %, wind direction: degree from 0 to 360 with 0 as North.
 Following are some examples of the input and output of the task, you can make
 slightly changes to the output to make it more natural and fluent, but keep the
 main information, concise and the ranking instructions in mind:
 Example Input 1:
 xxxxxxx
 Example Output 1-1:
 It’s the early morning of a day in January.
 The current weather is clear.
 The weather is expected to remain clear.
 The temperature is showing a mild drop.
 There is Light Breeze from NNW.
 The atmospheric shows Average Pressure.
 The humidity is very high.
 Example Output 1-2:
 It’s the early morning of a day in January.
 The current weather is clear.
 The weather will keep clear.
 The temperature is dropping mildly.
 There is Light Breeze from NNW.
 The atmospheric pressure is average.
 The air is very humid.
K.3 Channel Details
The meaning of each channel are as follows. We ask GPT4 to generate this document. There might be incorrectness. These sentences are then embedded as channel description embeddings used in the dataset.
  • p (mbar): Atmospheric pressure measured in millibars. It indicates the weight of the air above the point of measurement.
  • T (degC): Temperature at the point of observation, measured in degrees Celsius.
  • Tpot (K): Potential temperature, given in Kelvin. This is the temperature that a parcel of air would have if it were brought adiabatically to a standard reference pressure, often used to compare temperatures at different pressures in a thermodynamically consistent way.
  • Tdew (degC): Dew point temperature in degrees Celsius. It’s the temperature to which air must be cooled, at constant pressure and water vapor content, for saturation to occur. A lower dew point means dryer air.
  • rh (%): Relative humidity, expressed as a percentage. It measures the amount of moisture in the air relative to the maximum amount of moisture the air can hold at that temperature.
  • VPmax (mbar): Maximum vapor pressure, in millibars. It represents the maximum amount of moisture that the air can hold at a given temperature.
  • VPact (mbar): Actual vapor pressure, in millibars. It’s the current amount of water vapor present in the air.
  • VPdef (mbar): Vapor pressure deficit, in millibars. The difference between the maximum vapor pressure and the actual vapor pressure; it indicates how much more moisture the air can hold before saturation.
  • sh (g/kg): Specific humidity, the mass of water vapor in a given mass of air, including the water vapor. It’s measured in grams of water vapor per kilogram of air.
  • H2OC (mmol/mol): Water vapor concentration, expressed in millimoles of water per mole of air. It’s another way to quantify the amount of moisture in the air.
  • rho (g/m³): Air density, measured in grams per cubic meter. It indicates the mass of air in a given volume and varies with temperature, pressure, and moisture content.
  • wv (m/s): Wind velocity, the speed of the wind measured in meters per second.
  • max. wv (m/s): Maximum wind velocity observed in the given time period, measured in meters per second.
  • wd (deg): Wind direction, in degrees from true north. This indicates the direction from which the wind is coming.
  • rain (mm): Rainfall amount, measured in millimeters. It indicates how much rain has fallen during the observation period.
  • raining (s): Duration of rainfall, measured in seconds. It specifies how long it has rained during the observation period.
  • SWDR (W/m²): Shortwave Downward Radiation, the amount of solar radiation reaching the ground, measured in watts per square meter.
  • PAR (umol/m2̂/s): Photosynthetically Active Radiation, the amount of light available for photosynthesis, measured in micromoles of photons per square meter per second.
  • max. PAR (umol/m2̂/s): Maximum Photosynthetically Active Radiation observed in the given time period, indicating the peak light availability for photosynthesis.
  • Tlog (degC): Likely a logged temperature measurement in degrees Celsius. It could be a specific type of temperature measurement or recording method used in the dataset.
  • CO2 (ppm): Carbon dioxide concentration in the air, measured in parts per million. It’s a key greenhouse gas and indicator of air quality.
K.4 Dataset Structure
Weather-Captioned/
|-- caption_emb/   // Pre-embeddings
|   |-- somehash0.npy
|   |-- somehash1.npy
|-- caption_emb_large/
|   |-- somehash0.npy
|   |-- somehash1.npy
|   |-- somehash2.npy
|   |-- somehash3.npy
|-- utilities/    // utilities for spider, cleaning, embedding
|   |-- spider.ipynb
|   |-- data_clean.ipynb
|-- raw_data/   // raw weather report data
|   |-- daily_weather20xx.csv
|   |-- 6hourly_weather20xx.csv
|-- captioned_data/   // GPT captioned sentences
|   |-- weather_messages20xx.csv
|-- hashtable.json      // hash table for sentences and its hash
|-- hashtable_large.json
|-- weather_2014-18.parquet     // time series data
|-- weather_large.parquet
Appendix L Steam Details
L.1 Data source
We are not directly publishing this dataset because of the intellectual property restrictions.
For those who are interested in reproducing the dataset: We directly crawl all the event data on Steam News of each Game. We may release the script for this later.
As for the online players, we use a randomly picked publicly available database that provide the downloading of csv file containing online gamer number. We will not release any data or tools to get these data.
Appendix M Implementation Details and Hyper-Parameters
We train our model on single NVIDIA A800 GPU.
For electricity dataset, we directly report the result from the original paper. For weather dataset, we uses the exact set of hyper-parameter for the original weather datasets provided by each baseline model.
In most of the experiments, we simply use a patch length of 6 and stride of 3. For Toy dataset, we use patch length of 16 and stride of 8.
We follow the previous works, split all the dataset by 7:1:2 for training, validation and testing.
Except the performance on the Weather-Captioned Dataset, all other experiments are ran on the MiniLM Embedding.
Further detailed hyperparameter settings are provided in the training scripts in our codebase. We did not perform comprehensive hyper-parameter tuning because of the constraint of compute power. Thus, we may report a sub-optimal result of TGForecaster.