### v3

# Intervention-Aware Forecasting: Breaking Historical Limits from a System Perspective

Abstract
Traditional time series forecasting methods predominantly rely on historical data patterns, neglecting external interventions that significantly shape future dynamics. Through control-theoretic analysis, we show that the implicit "self-stimulation" assumption limits the accuracy of these forecasts. To overcome this limitation, we propose an Intervention-Aware Time Series Forecasting (IATSF) framework explicitly designed to incorporate external interventions. We particularly emphasize textual interventions due to their unique capability to represent qualitative or uncertain influences inadequately captured by conventional exogenous variables. We propose a leak-free benchmark composed of temporally synchronized textual intervention data across synthetic and real-world scenarios. To rigorously evaluate IATSF, we develop FIATS, a lightweight forecasting model that integrates textual interventions through Channel-Aware Adaptive Sensitivity Modeling (CASM) and Channel-Aware Parameter Sharing (CAPS) mechanisms, enabling the model to adjust its sensitivity to interventions and historical data in a channel-specific manner. Extensive empirical evaluations confirm that FIATS surpasses state-of-the-art methods, highlighting that forecasting improvements stem explicitly from modeling external interventions rather than increased model complexity alone.
1 Introduction
Time series forecasting (TSF) has witnessed significant progress, yet recent studies indicate diminishing returns: deep learning models [1, 2, 3] or even pretrained time series foundation models [4, 5, 6] now deliver only marginal performance gains over simple linear baselines [7, 8, 9].
This performance plateau arises primarily because traditional TSF methods rely exclusively on historical data, inherently adopting a problematic "self-stimulation" assumption—forecasting models depend solely on past observations while ignoring external interventions. In reality, time series often originate from dynamic systems that evolve not just from their previous state but also through external interventions. With a control-theoretic framework, we demonstrate that this modeling gap imposes an insurmountable barrier on forecasting accuracy. Recent studies have made preliminary yet promising attempts to incorporate both textual [10, 11, 12] and exogenous variables [13, 14] context for forecasting, though lacking rigorous theoretical grounding. Our analytical framework explicitly demonstrates that incorporating intervention-related context can lower forecasting error bounds.
Motivated by this insight, we propose Intervention-Aware Time Series Forecasting (IATSF), a novel forecasting paradigm that incorporates external interventions into conditional predictions. Since interventions may take diverse, often qualitative forms, we focus on textual data due to its ubiquity and ability to encode nuanced, non-quantifiable signals. By modeling interventions explicitly, IATSF reframes forecasting from correlation-based inference to dynamic system modeling, providing a principled framework for integrating textual context as supplementary intervention signals. This approach not only aligns forecasting with real-world system dynamics but also offers practical advantages in interpretability and adaptability.
Despite these theoretical advances, practical adoption remains challenging due to the lack of datasets and models that are compatible with intervention-aware forecasting. Existing multimodal time series forecasting (TSF) approaches often rely on large language models (LLMs) and datasets [15, 10] optimized for prompting rather than structured intervention modeling. Consequently, these datasets often have: (1) short horizons limiting meaningful intervention evaluation; (2) overly simplistic or ambiguous textual descriptions causing information leakage or irrelevance; and (3) poor temporal synchronization between textual and numerical data. To address these limitations and operationalize our theoretical insights, we introduce the Temporal-Synced IATSF benchmark, explicitly designed with leak-free textual interventions synchronized to extended, realistic forecasting horizons.
To demonstrate the effectiveness of intervention-aware forecasting, we propose FIATS (Forecaster for Intervention-Aware Time Series), a lightweight, LLM-free baseline model. FIATS uses semantically aligned textual embeddings and introduces a novel Channel-Aware Adaptive Sensitivity Modeling (CASM) mechanism guided by control theory. Additionally, a Causal Alignment Decoder with Channel-Aware Parameter Sharing (CAPS) explicitly aligns textual interventions with forecasting channels. Extensive experiments across synthetic, physics-based, and market datasets show that FIATS consistently outperforms state-of-the-art methods, with ablation studies confirming that performance improvements stem from explicit intervention modeling rather than model complexity.
In summary, our key contributions are:
  • A control-theoretic analysis reveals intrinsic forecasting barriers caused by the "self-stimulation" assumption (sole reliance on history) and shows intervention-aware modeling reduces error bounds.
  • Building on this analysis, we introduce IATSF, a paradigm that models time series with external interventions, bridging the gap between traditional TSF and real-world dynamic systems.
  • We operationalize IATSF with the Temporal-Synced IATSF benchmark and a LLM-free FIATS model, whose performance gains are shown to stem from principled intervention modeling, not architectural complexity.
2 Background and Motivation: TSF from System Analysis Perspective
Time series data are typically measurements of real-world dynamic systems whose behaviors are continually shaped by external events. However, conventional datasets and forecasting methods often rely exclusively on historical measurements, neglecting these influential external factors. For instance, the widely-used ETT dataset [16] records power load and oil temperature from electric transformers—both significantly impacted by external events, including human activities and environmental conditions. Traditional approaches incorporating numeric exogenous variables, such as ARIMAX [17], have advanced forecasting capabilities by explicitly including external numeric inputs. Nevertheless, these methods fall short when dealing with qualitative, uncertain external factors frequently represented in textual form—such as event descriptions, news reports, or expert narratives. Recent works have attempted to bridge this gap by incorporating textual contextual information to improve forecasting accuracy [10, 11, 12], though a rigorous theoretical justification for their effectiveness remains lacking.
To systematically address this qualitative gap, we formally identify and analyze the intrinsic limitations of ignoring qualitative external interventions from a dynamical systems perspective [^1] All proofs and discussion are provided in Appendix B, unless otherwise specified..
2.1 Time Series Are Observation of Real-World Dynamic Systems
Consider a general dynamical system characterized by hidden states $Z\in\mathbb{R}^{m}$, evolving based on historical states and independent external interventions [18, 19, 20]:
$Z_{f}=F(Z_{h},U_{t}),\quad X=O(Z)$ (1)
where $F$ represents the true system dynamics, $U_{t}$ denotes time-varying independent external interventions, $O$ represents observation, $X$ for the the observed signal. For analytical clarity, we assume full observability, i.e. $X=Z$. We also discuss a simple linear system case $X_{f}=AX_{h}+BU_{t}$, where $A$ governs self-stimulated state transitions and $B$ encodes intervention sensitivity. Standard forecasting datasets $\mathcal{D}=\{(X_{h}^{(i)},X_{f}^{(i)})\}_{i=1}^{N}$ are generated through sliding window on the observed signals, where $X_{h},X_{f}$ stand for look-back window and forecasting horizon segment accordingly.
2.2 The Implicit Self-Stimulation Assumption in TSF
Traditional forecasting adopts a self-stimulation paradigm where models $f_{\theta}$ attempt to approximate system dynamics using only historical observations:
$f_{\theta}^{*}=\arg\min_{\theta}\mathbb{E}\left[\|\epsilon\|^{2}\right]=\arg\min_{\theta}\mathbb{E}\left[\|F(X_{h},U_{t})-f_{\theta}(X_{h})\|^{2}\right]$ (2)
The critical limitation stems from implicitly treating unobserved interventions as hidden random variables $U_{t}\sim\mathcal{P}_{U}$. This induces an irreducible forecasting error, as formalized by our first proposition:
Proposition 2.1 (Self-Stimulation Error Bound).
For any self-stimulated model $f_{\theta}$, it converges to predicting conditional expectation $F^{*}(X_{h},\mu)\triangleq\mathbb{E}_{U}[F(X_{h},U)]$, the prediction error covariance satisfies:
$\text{Cov}(\epsilon)\succeq\mathbb{E}_{X_{h}}\left[\nabla_{U}F\Sigma(\nabla_{U}F)^{\top}\right]$ (3)
where $\mu=\mathbb{E}(U_{t}),\quad\Sigma=\text{Cov}(U_{t})$. For linear systems, this falls back to:
$\text{Cov}(\epsilon)\succeq B\Sigma B^{\top}$ (4)
Proposition 2.1 reveals two fundamental limitations: 1) Self-stimulated models converge to predicting conditional expectations, rather than true dynamics, explaining prevalent averaging effects in practice as shown in Fig. 1, and 2) An irreducible error floor exists due to intervention stochasticity. This establishes a theoretical performance ceiling for conventional TSF approaches.
Figure 1: The real system runs under various interventions. The Intervention-Aware method can effectively approximate the real system according to the dataset while traditional self-stimulated method can only approximate a average scenario with persistent error, lead to bad or even collapse result. The right panel shows visualization result of a frequency modulated system which is very sensitive to the intervention, i.e. large $\nabla_{U}F$.
3 IATSF: Intervention-Aware Time Series Forecasting
3.1 Task Formulation
We propose Intervention-Aware Time Series Forecasting (IATSF) to overcome the self-stimulation limitation. The key innovation lies in explicit intervention modeling:
$f_{\theta}^{*}=\arg\min_{\theta}\mathbb{E}\left[\|F(X_{h},U_{t})-f_{\theta}(X_{h},U_{t})\|^{2}\right]$ (5)
where $U_{t}$ represents measurable interventions. This paradigm enables breaking the error bound in proposition 2.1 through intervention-aware learning, as detailed in Fig. 1.
As shown above, instead of assuming the external intervention stays the same in the TSF, IATSF aims to predict a conditioned future with the observed or predicted intervention even though it is not fully observed or precise. The error reduction mechanism is formalized through our second proposition:
Proposition 3.1 (Partial Intervention Efficacy).
For a system with $p$ independent interventions $U_{t}=\sum_{i=1}^{p}U_{t}^{i}$, incorporating any known intervention $U_{t}^{j}$ reduces the error covariance by:
$\Delta\text{Cov}(\epsilon)=\nabla_{U_{j}}F\Sigma_{j}(\nabla_{U_{j}}F)^{\top}$ (6)
For linear systems, this reduces the lower bound by $B_{j}\Sigma_{j}B_{j}^{\top}$.
Proposition 3.1 demonstrates that any measurable intervention information reduces forecasting uncertainty, even with incomplete intervention knowledge. This motivates our key insight: textual descriptions of interventions provide viable information for uncertainty reduction, despite non-numeric formats.
3.2 Language as an Intervention Modality
Incorporating exogenous variables is a common approach [13, 14], but it typically requires numerical time series or one-hot encoded inputs sampled at the same rate as the target series—even when the actual interventions are sparse. This limits flexibility, especially when new events occur. In real-world settings, many impactful factors—such as weather anomalies, geopolitical shifts, or human decisions—are hard to quantify but still essential for accurate forecasting. To address this, we propose modeling interventions using linguistic descriptors, which naturally capture compositional and relational semantics through lexical encoding. This allows for expressive representations of complex events (e.g., "simultaneous port strikes and agricultural subsidies") without incurring combinatorial overhead. This design offers several key advantages:
Expert Knowledge Integration: Textual interfaces facilitate the direct inclusion of domain-specific expertise via natural language specifications (e.g., "anticipated regulatory changes will suppress industrial output"). This makes it easier to incorporate human input or LLM-driven forecasting through linguistic conditioning of interventions.
Generalizability: Textual representations provide flexibility across various contexts, allowing models to generalize more effectively to new or unseen intervention scenarios. The use of natural language reduces reliance on rigid, pre-encoded numerical data, enabling better adaptability to diverse situations.
Cross-Modal Causal Alignment: By embedding both linguistic intervention descriptors and their temporal effects in a shared space, neural architectures can learn latent mappings that align interventions with their causal impacts on the system.
4 IATSF Benchmark Datasets
4.1 Leak-Free Dataset Design
The IATSF benchmark is explicitly constructed to be leak-free, adhering to the principle that models must not access future system states. To enforce this, we only include independently evolving interventions—external causal factors that influence the system but are not themselves outcomes of it. Including variables that directly describe or summarize the time series trajectory (as in [15, 3]) would violate this principle by introducing future state information; see Appendix O for further discussion.
Since system responses to interventions often occur much faster than the sampling interval (e.g., photovoltaic panels react to sunlight in milliseconds), we assume interventions take effect instantaneously and denote the up-to-date intervention as $U_{f}$. However, in real-world deployment, such ground-truth interventions are unavailable at prediction time. Therefore, we restrict the intervention input to three categories: (1) Known information, such as holidays or other common knowledge; (2) Predictions of $U_{f}$ derived from sources with expert knowledge, such as weather reports; and (3) Hypothetical or controlled events, which allow the IATSF models to simulate "what-if" scenarios during decision-making. Evaluation strategies accounting for prediction errors in interventions are detailed in Appendix B.3.
4.2 Brief Datasets Introduction
Each instance in IATSF is defined as $\mathcal{D}=\left\{((X_{h}^{(i)},U_{f}^{(i)},D),X_{f}^{(i)})\right\}_{i=1}^{N}$, comprising historical time series $X_{h}$, future-aligned interventions $U_{f}$, with $D$ denoting channel descriptors and future values $X_{f}$ as ground truth. $U_{f}$ may contain multiple temporally-aligned interventions, each independently observed alongside the time series. The datasets contains multiple channels or instances, each with distinct distributions.
As mentioned earlier, the key challenge is to identify time-synced interventions that are independently observed alongside the time series. This makes it impractical to apply traditional datasets, such as ETT [16], which lack the necessary contextual information. To address this, we have designed four initial IATSF benchmark datasets across four domains: synthetic controlled systems, physics, building management, and market analysis. Each dataset includes temporally aligned interventions and time series data, making them suitable for IATSF validation. Please check Appendix P for detail.
Frequency Modulated Toy Dataset
A simple intervention-aware system with sinusoidal wave segments and varying frequencies under the influence of interventions. Textual descriptions precede each change point, providing a clear context for upcoming alterations. With full intervention observation, this dataset has a theoretical error lower bound of 0.
Electricity Utility Dataset
Based on a widely used dataset for office building appliance usage [16], this dataset incorporates daily patterns affected by workdays. We enhance it with textual information (e.g., day type, public holidays), using channel names as descriptors. This dataset allows us to explore the impact of minimal textual data on prediction accuracy and compare it to traditional models.
Atmospheric Physics Dataset
Sourced from a research initiative monitoring fine-grained atmospheric signals. In addition to commonly reported variables such as temperature and humidity, it includes measurements like Dew Point and Short-Wave Downward Radiation (SWDR), which are closely linked to weather interventions but are excluded from general weather reports. This dataset observes an ideal system for studying IATSF, as it provides clean, direct, and highly correlated intervention effects. We use open-source weather report APIs to prepare textual interventions, thereby avoiding direct access to the time series data. By incorporating limited system-level predictions, this dataset demonstrates how expert knowledge and external information can improve the accuracy of fine-grained time series forecasting. It also includes multiple channels with distinct distributions and intervention responses, presenting challenges in channel-specific behavior modeling.
Game Active User Dataset (GAUD)
A key application of IATSF is to model the impact of business decision to the market. This dataset tracks daily active users for 90 games on an online platform, with developer, category, and update logs as interventions. It helps evaluate the model’s ability to capture market response to human interventions. The dataset is highly random with complex nonlinearities and includes short time series, allowing for evaluation in cold-start or zero-shot scenarios.
5 FIATS: A Simple System-Aware Baseline Model for IATSF
While recent studies [11, 10, 15, 12, 21] explore text-informed forecasting using the reasoning capabilities of large language models (LLMs), these approaches are limited to short sequences and simplistic interventions. They often suffer from high variance, low interpretability, and significant computational and token overhead. To address these limitations and rigorously validate the IATSF task, we introduce FIATS—the first LLM-free, numerical-based forecaster designed for intervention-aware time series. As illustrated in Fig. 2, FIATS combines a patch-based time series encoder [1] with novel text-embedding-based [22, 23] intervention semantic encoder and decoder. The novelty is as follows:
Temporal-Synced Intervention
Real-world systems often respond rapidly to interventions, necessitating temporal alignment between text and time series data. FIATS addresses this by synchronizing each time series patch with the last intervention observed, e.g. for patch start from 10:15, sync with the last timestep with intervention update of 10:00. This ensures the model uses only leak-free, contemporaneous interventions when forecasting subsequent patches, preventing future information leakage while maintaining temporal relevance.
Channel-aware Adaptive Sensitivity Modeling (CASM)
In FIATS, we reframe the attention mechanism through a control-theory perspective to explicitly model intervention sensitivity—a novel approach within attention-based architectures. Starting from linear systems where time series are observed by $X_{f}=CZ_{f}=CAZ_{h}+CBU_{f}$, channel-specific sensitivity to interventions is governed by $\frac{dx_{f}^{i}}{dU_{f}}=c^{i}B$. This indicates that each channel responds differently to external interventions. The error analysis is discussed in Appendix B.4. To capture this without introducing excessive parameters, we reconceptualize cross-attention as a Channel-aware Adaptive Sensitivity Modeling Block, as shown in the right panel of the Fig. 2, specifically:
- Query as Channel-wise Sensitivity $\tilde{C}=Desc\cdot W_{Q}$: Channel descriptions $Desc\in\mathbb{R}^{CN\times D}$ are served as query ($CN$ as channel number). The query projection explicitly learns how textual channel features (e.g., "atmospheric pressure") influence intervention sensitivity for each channel. This allows the model to adjust how interventions are perceived based on channel-specific characteristics.
- Key as Intervention Filter $\tilde{B}_{U_{f}}=(News\cdot W_{K})^{\top}$: The key projection maps temporal-synced news embeddings $News\in\mathbb{R}^{M\times D}$ to a system sensitivity matrix ($M$ as news number), allowing the model to filter out irrelevant interventions (e.g., excluding "tech stock news" when forecasting atmospheric physics). This ensures that only pertinent interventions are considered for each system.
- Value as Intervention Translator $\tilde{U_{f}}=News\cdot W_{V}$: Value projection learns to maps news text embedding to $\tilde{U_{f}}$, the latent space of actionable intervention effects.
Figure 2: Architecture of FIATS. FIATS integrates three inputs: time series data from a look-back window, temporal-synced news embeddings, and channel description embeddings. The intervention encoder employs CASM blocks in a residual connection along with multiple self-attention layers to enhance feature extraction. The CAPS causal alignment decoder projects the historical time series embeddings into the future, guided by channel-aware, time-synced interventions. A token-wise decoder is used to prevent overfitting in the final linear layer, as discussed in [24].
The above analysis show that the attention mechanism can effectively generate the channel-aware intervention $U_{f}^{c}$. This design allows identical interventions to differentially impact channels based on their descriptions. Unlike static sensitivity coefficients found in classical systems, this formulation maintains the nonlinear characteristics provided by the transformer block, allowing for greater learning flexibility to approximate complex nonlinear system. Additionally, it aligns well with the theoretical framework, making the model more interpretable. The attention map produced by the CASM layer directly reveals the sensitivity of each channel to various interventions, providing clear insights into how interventions impact different channels based on their specific descriptions.
Channel-Aware Parameter Sharing (CAPS)
While CASM addresses heterogeneous intervention responses, channels also exhibit inherent differences in their temporal patterns – a critical factor neglected by conventional parameter sharing. Previous shared models approximate all channels with a same set of parameters introducing persistent errors $\epsilon_{i}=o_{i}(Z)-\frac{1}{k}\sum_{j=1}^{k}o_{j}(Z)$ where $o_{i}$ for real system channel-specific dynamics.
To mitigate this issue, FIATS introduces a lightweight channel-aware decoding mechanism. All channels are first encoded into a shared latent space $\tilde{Z}$ by a unified time-series encoder. Then, a channel-conditioned decoder is used to adaptively project this latent representation into a channel-aware space, conditioned by the channel-specific time-synced intervention embeddings $U_{f}^{c}$. decoder approximates channel-specific adjustments by modulating the shared latent space through cross-attention $Attention(Q=U^{c}_{t},K,V=\tilde{Z})$ to simulate such nonlinear projection. To avoid future information leakage, we apply causal attention mask here. We will omit the analysis.
This design introduces minimal overhead while enabling the model to account for channel heterogeneity in a flexible, data-driven manner. Additionally, the attention maps produced by the channel-aware decoder are interpretable: they reveal how each channel selectively attends to historical time series data under different interventions. We provide visualizations and further analysis of these attention patterns in the following session.
6 Experiments
Baseline Models
FIATS is benchmarked against several state-of-the-art (SOTA) methods. These include linear-based models [7, 8] as , transformer-based models [1, 2], and fine-tuned LLM-based multimodal method [3]. Additionally, we compare pretrained time series "foundation models" [4, 6, 5]. This selection covers a range of approaches, including self-stimulated linear and nonlinear models, data-specific and pretrained models, LLM-based cross-modal models.
6.1 Evaluation on FM Toy Dataset
The FM Toy Dataset is generated using a fully-controlled frequency modulation system with a single intervention factor. This dataset has a theoretical error lower bound of 0, providing an ideal environment to analyze model performance.
Statistical Results
Table 1 shows that FIATS achieves near-zero error, closely aligning with the theoretical lower bound, while other TSF methods, including pre-trained models, exhibit considerably higher error even with simple sinusoidal data. This confirms that the primary bottleneck is the self-stimulation assumption, not the quantity or variety of data.
Notably, linear-based models perform poorly in this intervention-sensitive context due to their limited parameters and linearity, resulting in collapsed predictions. In contrast, PatchTST, the best-performing self-stimulated model, demonstrates robustness in handling nonlinearity. As seen in Fig. 1, PatchTST captures periodicity, but with diminishing amplitude over time. This aligns with Proposition 2.1, which states that as the prediction horizon extends, the system’s behavior is more likely to be influenced by new interventions. As a result, self-stimulated models tend to predict conditional expectations in far future, leading to diminished amplitude. FIATS, by contrast, outperforms these models at longer horizons, as it incorporates the increasing influence of interventions while the impact of historical data fades.
Table 1: Forecasting result in MSE, comparing the intervention-aware FIATS against various TSF methods. The best result is highlighted in bold and the second best is highlighted in underscore.
Dataset
Pred.
Len.
FIATS FITS DLinear PatchTST iTrans. Chronos-L MOIRAI-L Time-MoE-U TimeLLM
14 0.003 0.282 0.151 0.006 0.136 0.012 0.013 0.012 0.231
28 0.008 0.692 0.297 0.029 0.295 0.047 0.062 0.035 0.382
60 0.020 0.909 0.442 0.075 0.494 0.129 0.133 0.107 0.551
FM Toy 120 0.027 0.883 0.632 0.168 0.747 0.374 0.385 0.295 0.788
96 0.124 0.134 0.140 0.130 0.148 0.154 0.152 0.149 0.131
192 0.144 0.149 0.153 0.149 0.162 0.177 0.171 0.168 0.152
336 0.158 0.165 0.169 0.166 0.178 0.197 0.192 0.183 0.160
Electricity Utility 720 0.190 0.203 0.204 0.210 0.225 0.242 0.236 0.229 0.192
96 0.182 0.248 0.294 0.252 0.267 0.293 0.299 0.258 0.294
192 0.205 0.297 0.340 0.304 0.327 0.357 0.356 0.318 0.342
336 0.235 0.354 0.393 0.364 0.404 0.448 0.457 0.413 0.393
Atmospheric Physics 2014-19 720 0.281 0.430 0.456 0.439 0.495 0.512 0.532 0.508 0.461
96 0.410 0.436 0.487 0.464 0.456 0.447 0.453 0.437 -
192 0.438 0.524 0.568 0.567 0.578 0.552 0.557 0.542 -
336 0.455 0.601 0.644 0.644 0.698 0.685 0.673 0.647 -
Atmospheric Physics 2014-24 720 0.497 0.692 0.725 0.745 0.832 0.754 0.765 0.734 -
6.2 Evaluation on Real World Dynamic System
Statistical Results
On the Electricity dataset, FIATS demonstrates SOTA performance, particularly effective at shorter forecasting horizons using minimal textual cues. It is interestingly to see that as the prediction length gets longer, the irregularly appeared holiday events tend to contribute less to the overall loss since the loss is averaged out. The TimeLLM with large-language-model backbone does show some capability to perform forecast according to such simple intervention information and obvious causal correlation.
Table 2: A selection of channel-wise performance on Atmos. Phy. 2014-19 dataset in MSE.
Channel FIATS FITS DLinear PatchTST iTrans. IMP.
p (mbar) 0.136 0.863 0.823 0.930 1.032 83.43%
Tpot (K) 0.182 0.316 0.352 0.322 0.353 42.18%
VPdef (mbar) 0.283 0.638 0.696 0.674 0.803 55.59%
rho (g/m³) 0.192 0.390 0.411 0.418 0.453 50.73%
raining (s) 0.790 0.873 0.937 0.859 0.994 8.04%
SWDR (W/m²) 0.182 0.308 0.385 0.296 0.377 38.39%
As shown in Table 1, FIATS consistently outperforms all baselines on the Atmospheric Physics dataset. These results underscore that incorporating external intervention information directly addresses the information insufficiency in conventional TSF models. In contrast, pretrained TSF models, such as Chronos-L and MOIRAI-L, despite being trained on larger datasets, still underperform FIATS—highlighting that scaling data alone cannot compensate for the fundamental limitations imposed by the self-stimulation assumption and the absence of intervention modeling.
Table 2 further breaks down the results by channel on the Atmospheric Physics dataset. FIATS shows substantial performance gains on variables such as pressure ($p$), air density ($\rho$), and vapor pressure (VPdef)—channels not directly referenced in the collected weather report. This demonstrates FIATS’s ability to perform channel-aware modeling and infer latent causal relationships between interventions and time series patterns, even when the correlation is indirect or not explicitly observed. The full breakdown performance shown in Appendix G.
6.3 Evaluation on Market System
We evaluate FIATS on the GAUD dataset to test its ability to handle real-world, intervention-driven market dynamics. Each time series tracks daily active users of a game over a 60-day input window and a 14-day forecast horizon. Due to large developer variability and temporal shifts, traditional TSF models struggle to generalize across games. FIATS addresses this by incorporating textual data, enabling intervention-aware forecasting.
Figure 3: Performance improvement with respect to the PatchTST on each time series in GAUD.
As shown in Fig. 3, FIATS consistently outperforms PatchTST, achieving an average improvement of 12.6% and ranking first on 59.6% of the games. In some cases, it boosts accuracy by up to 50–80%. Notably, for games released after 2021, where time series are short and cold-start issues arise, FIATS shows clear advantages. Its ability to generalize from textual interventions allows it to maintain performance where self-stimulated models like PatchTST fail to converge. Compared to TimeLLM, which depends on prompt templates, FIATS leverages raw textual semantics more effectively, resulting in broader applicability and better accuracy. Full results are provided in Appendix M.
6.4 Case Study & Ablation Study
Case Study: Visualization and Controllability Test
Fig. 4 visualizes three representative channels from the Atmospheric Physics dataset (full results in Appendix H). The first channel, atmospheric pressure ($p$), is sensitive to regional climate shifts but lacks strong short-term historical correlation. Its slow, subtle changes are challenging for traditional TSF models. PatchTST fails to capture these dynamics, defaulting to a flat prediction, while FIATS successfully models the trend by conditioning on relevant interventions.
Figure 4: Visualization of three channels on the 15,000th test sample of the Atmos. Phy. 2014-19 dataset. Blue indicates ground truth, Red shows FIATS, Green represents PatchTST, and Orange denotes FIATS with swapped interventions on the second and fourth forecast days. The CAPS causal alignment decoder exhibits distinct attention patterns across channels.
The second channel, rainfall duration (in seconds per 10 minutes), is sparse and lacks periodicity. PatchTST outputs near-zero values—its conditional expectation under uncertainty—while FIATS adjusts its predictions based on available intervention signals. It correctly forecasts the first rainfall event but misses the second due to misaligned or absent external information, reflecting dependence on accurate intervention input.
The third channel, solar radiation (SWDR), is not explicitly mentioned in the intervention but is indirectly influenced. FIATS captures its phase and amplitude accurately, thanks to the CASM design that enables cross-channel sensitivity modeling. PatchTST, by comparison, produces generic, misaligned waveforms.
A controllability test, Swapping interventions on the second and fourth forecast days confirms FlATS’s responsiveness. It updates predictions accordingly—forecasting rain on the fourth and clear skies on the second—demonstrating its ability to adapt to changes in external interventions in a causally aligned manner.
Case Study: Attention Map for Interpretability
: The CASM block analysis in Fig. 5 shows how the model focuses on different temporal features across layers. In the first layer, attention centers on the first sentence, providing temporal context for daily and annual periodicity. The second layer shifts attention to channel-specific signals, particularly the sixth sentence describing atmospheric pressure, reflecting the model’s sensitivity to channel-specific patterns and interventions. By the third layer, attention diversifies, focusing on relevant intervention aspects for each channel.
(a) CASM Layer1
(b) CASM Layer2
(c) CASM Layer3
Figure 5: Attention map of the CASM on the 15000th test sample of Atmos. Phy. 2014-19 dataset. We use three cross attention block in residual connection. The horizontal axis stands for channels and vertical stands for the 7 sentences of the weather report summary.
The CAPS causal alignment decoder, shown in Fig. 4, demonstrates distinct attention patterns across channels, highlighting the model’s ability to align time series data with textual interventions. Channels associated with periodic variables like SWDR exhibit clear periodicity in attention maps, indicating effective capture of cyclical patterns. Rainfall channel highlights historical rainfall, showcasing the model’s sensitivity to key moments. This adaptability is driven by CASM, enabling the model to tailor its attention based on each channel’s unique characteristics. Analysis on such attention map may further reveal the underlying causal correlation about how a time series or a system reacts to certain external intervention for future work. Full analysis see Appendix K.
Table 3: Ablation result on Atmos. Phy. 2014-19 in MSE.
Pred.
Len.
Openai
512
MiniLM mpnet
Zero
Desc.
Zero
News
96 0.182 0.186 0.196 0.209 0.249
192 0.205 0.214 0.216 0.260 0.302
336 0.235 0.232 0.251 0.302 0.359
720 0.281 0.272 0.291 0.356 0.432
Ablation: Effectiveness & Robustness
We evaluate FIATS’s performance across different text embedding spaces by switching the embedding model. As shown in Table 3, the results reveal minimal performance variation, demonstrating the generalizability of the FIATS architecture.
Figure 6: Loss under various noise levels. Blue line for horizon 96 and Orange line for horizon 192.
Next, we add random noise to the news embeddings to simulate imperfect intervention inputs. In Fig. 6, minor noise that does not alter sentence semantics—correspond to slightly changing wording while preserving the overall meaning—has minimal impact on model performance, showing the semantic robustness of FIATS. However, as noise levels increase, the forecasting performance progressively degrades. This observation supports Proposition 3.1, highlighting that forecasting accuracy depends on the accuracy and coverage of the observed intervention relative to the true intervention. When intervention input is entirely randomized, FIATS’ performance deteriorates to match PatchTST, indicating that the model ignores meaningless intervention signals. Additionally, comparisons between the 192-step and 96-step forecast horizons show that longer forecasts are more sensitive to intervention noise, consistent with previous observations.
Finally, we mask the news and description with zero tensors to directly assess their contribution. In Table 3, removing news embeddings reduces performance to levels similar to randomized intervention input, underscoring that without meaningful interventions, the model behaves as a purely historical-data-driven (self-stimulated) model. Eliminating channel descriptions significantly worsens forecasting performance, demonstrating their critical role in accurately modeling channel-specific sensitivities. Additional ablation analyses related to causal alignment can be found in Appendix J.
7 Conclusion, Limitation & Future Work
This paper presents Intervention-Aware Time Series Forecasting (IATSF), leveraging a control-theoretic framework to address errors from the self-stimulation assumption and improve forecasting accuracy through intervention modeling. We demonstrate the effectiveness of IATSF using the Temporal-Synced IATSF benchmark and the FIATS model, which outperforms state-of-the-art methods, including those based on large language models. Our findings emphasize that intervention-aware modeling, rather than simply increasing model complexity, is crucial for enhancing forecasting performance. While FIATS shows some capability in noise tolerance and generalization, challenges persist in modeling complex chaotic systems, where interventions may not have immediate effects and varying credibility of news sources or temporal misalignment could lead to inaccurate intervention observations. Overcoming these challenges will require more advanced models, potentially benefiting from pretraining techniques. These areas will be explored in future research. Additionally, the analysis framework can inspire further exploration, such as modeling multichannel correlation.
Appendix A Some Notation Used in Paper
Table 4: Summary of Some Notations Used in the Paper
Symbol Description
$F$ Real System dynamics function
$\theta$ Parameters of the forecasting model
$f_{\theta}$ Forecasting function with parameters $\theta$
$O$ Observation function
$Z$ Hidden states of the system
$X$ Observed signal from the system
$X_{f}$ Future time Series Segment
$X_{h}$ Historical time Series Segment
$\hat{X}_{f}$ Forecasted future time series
$U_{t}$ Time varying external intervention
$U_{f}$ External intervention for the future segment
$\Sigma$ Covariance of the interventions $U_{t}$
$\mu$ Mean of the intervention distribution
$\mathcal{D}$ Set of time series samples
$W_{Q},W_{K},W_{V}$ Weights for Query, Key, and Value in the attention mechanism
Appendix B Proof and Discussion
In this section, we give proof of the two proposition mentioned in the paper. We also discuss the error introduced by the external intervention forecaster, weight sharing and incomplete observation.
B.1 Proposition 2.1: Error Bound Introduced by Self-Stimulation
B.1.1 Most Simple Case: Linear System, Linear Model
Proof.
Consider a linear system with unobserved interventions $U$:
$X_{f}=AX_{h}+BU,\quad U\sim\mathcal{P}_{U}\ \text{(i.i.d.)},\ \mathbb{E}[U]=\mu,\ \text{Cov}(U)=\Sigma,$ (7)
where $X_{h}$ represents historical states and $X_{f}$ represents future states. We aim to estimate $X_{f}$ using a self-stimulated linear model:
$\hat{X}_{f}=CX_{h}+d,$ (8)
where $C$ and $d$ are parameters to be estimated via least squares by minimizing the loss:
$\mathcal{L}(C,d)=\mathbb{E}\left[\|X_{f}-(CX_{h}+d)\|^{2}\right].$ (9)
To find the optimal parameters $C^{*}$ and $d^{*}$, we take derivatives with respect to $d$ and $C$ and set them to zero.
Taking the derivative with respect to $d$:
$\frac{\partial\mathcal{L}}{\partial d}=\mathbb{E}\left[-2(X_{f}-CX_{h}-d)\right]=0$
$\mathbb{E}[d]=\mathbb{E}[X_{f}-CX_{h}]$
So, for the optimal $C^{*}$, $d^{*}$ is:
$d^{*}=\mathbb{E}[X_{f}]-C^{*}\mathbb{E}[X_{h}].$ (10)
Substituting $X_{f}=AX_{h}+BU$:
$d^{*}=\mathbb{E}[AX_{h}+BU]-C^{*}\mathbb{E}[X_{h}]=A\mathbb{E}[X_{h}]+B\mathbb{E}[U]-C^{*}\mathbb{E}[X_{h}]$
$d^{*}=(A-C^{*})\mathbb{E}[X_{h}]+B\mu.$
Next, taking the (Fréchet) derivative with respect to $C$ (or considering element-wise derivatives $\frac{\partial\mathcal{L}}{\partial C_{ij}}$), we set $\nabla_{C}\mathcal{L}(C,d^{*})=0$:
$\mathbb{E}\left[\nabla_{C}\|X_{f}-CX_{h}-d^{*}\|^{2}\right]=0$
$\mathbb{E}\left[-2(X_{f}-CX_{h}-d^{*})X_{h}^{\top}\right]=0$
$\mathbb{E}\left[(X_{f}-CX_{h}-d^{*})X_{h}^{\top}\right]=0$
Substituting $X_{f}=AX_{h}+BU$ and $d^{*}=(A-C)\mathbb{E}[X_{h}]+B\mu$:
$\mathbb{E}\left[(AX_{h}+BU-CX_{h}-((A-C)\mathbb{E}[X_{h}]+B\mu))X_{h}^{\top}\right]=0$
$\mathbb{E}\left[((A-C)X_{h}+B(U-\mu)-(A-C)\mathbb{E}[X_{h}])X_{h}^{\top}\right]=0$
$\mathbb{E}\left[(A-C)(X_{h}-\mathbb{E}[X_{h}])X_{h}^{\top}\right]+\mathbb{E}\left[B(U-\mu)X_{h}^{\top}\right]=0$
Assuming $X_{h}$ and $U$ are independent (or at least $U-\mu$ is uncorrelated with $X_{h}$), $\mathbb{E}[(U-\mu)X_{h}^{\top}]=\mathbb{E}[U-\mu]\mathbb{E}[X_{h}^{\top}]=0\cdot\mathbb{E}[X_{h}^{\top}]=0$.
$(A-C)\mathbb{E}\left[(X_{h}-\mathbb{E}[X_{h}])X_{h}^{\top}\right]=0$
$(A-C)\left(\mathbb{E}[X_{h}X_{h}^{\top}]-\mathbb{E}[X_{h}]\mathbb{E}[X_{h}^{\top}]\right)=0$
$(A-C)\text{Cov}(X_{h})=0.$
If $\text{Cov}(X_{h})$ is invertible (i.e., its columns are linearly independent and it has full rank), then we must have $A-C=0$, which implies:
$C^{*}=A.$ (11)
Substituting $C^{*}=A$ back into the expression for $d^{*}$:
$d^{*}=(A-A)\mathbb{E}[X_{h}]+B\mu=B\mu.$ (12)
So the optimal parameters are:
$C^{*}=A,\quad d^{*}=B\mu.$ (13)
The prediction error is given by:
$\epsilon=X_{f}-\hat{X}_{f}=(AX_{h}+BU)-(C^{*}X_{h}+d^{*})=(A-C^{*})X_{h}+BU-d^{*}.$ (14)
Substituting the optimal parameters $C^{*}=A$ and $d^{*}=B\mu$:
$\epsilon=(A-A)X_{h}+BU-B\mu=B(U-\mu).$ (15)
The mean of the error is $\mathbb{E}[\epsilon]=\mathbb{E}[B(U-\mu)]=B(\mathbb{E}[U]-\mu)=B(\mu-\mu)=0$. The covariance of the error is:
$\text{Cov}(\epsilon)=\mathbb{E}\left[\epsilon\epsilon^{\top}\right]=\mathbb{E}\left[(B(U-\mu))(B(U-\mu))^{\top}\right]$ (16)
$\text{Cov}(\epsilon)=\mathbb{E}\left[B(U-\mu)(U-\mu)^{\top}B^{\top}\right]=B\mathbb{E}\left[(U-\mu)(U-\mu)^{\top}\right]B^{\top}$
Since $\text{Cov}(U)=\Sigma=\mathbb{E}\left[(U-\mu)(U-\mu)^{\top}\right]$,
$\text{Cov}(\epsilon)=B\Sigma B^{\top}.$ (17)
This covariance represents the irreducible error floor caused by the unobserved intervention $U$.
Let $F(X_{h},U)=AX_{h}+BU$ be the true underlying system for $X_{f}$. The gradient of $F$ with respect to $U$ is $\nabla_{U}F=B$. Thus, even with optimal parameters, the error covariance satisfies:
$\mathbb{E}[\epsilon\epsilon^{\top}]=B\Sigma B^{\top}=(\nabla_{U}F)\Sigma(\nabla_{U}F)^{\top}.$ (18)
If we consider a general form of a lower bound related to the influence of $U$, such as one involving an expectation over $X_{h}$, $\mathbb{E}_{X_{h}}\left[(\nabla_{U}F)\Sigma(\nabla_{U}F)^{\top}\right]$, in this linear case it simplifies directly to $B\Sigma B^{\top}$ because $\nabla_{U}F=B$ does not depend on $X_{h}$. Therefore:
$\mathbb{E}[\epsilon\epsilon^{\top}]\succeq\mathbb{E}_{X_{h}}\left[(\nabla_{U}F)\Sigma(\nabla_{U}F)^{\top}\right]$ (19)
where $\succeq$ denotes positive semi-definiteness (Löwner order). In this specific linear case, this holds with equality: $\mathbb{E}[\epsilon\epsilon^{\top}]=B\Sigma B^{\top}$. This lower bound arises from the unobserved intervention $U$. ∎
B.1.2 A Step Further: Linear System, Nonlinear Model
Proof.
Consider the same linear system with unobserved interventions $U$:
$X_{f}=AX_{h}+BU,\quad U\sim\mathcal{P}_{U}\ \text{(i.i.d.)},\ \mathbb{E}[U]=\mu,\ \text{Cov}(U)=\Sigma.$ (20)
We now use a nonlinear self-stimulated model (e.g., an arbitrary machine learning model) for prediction:
$\hat{X}_{f}=f(X_{h}).$ (21)
The optimal self-stimulated model $f_{\text{opt}}(X_{h})$ that minimizes the Mean Squared Error (MSE) is the conditional expectation of $X_{f}$ given $X_{h}$:
$f_{\text{opt}}(X_{h})=\mathbb{E}[X_{f}\mid X_{h}]=\mathbb{E}[AX_{h}+BU\mid X_{h}].$
Assuming $U$ is independent of $X_{h}$ ($U\perp X_{h}$), then $\mathbb{E}[U\mid X_{h}]=\mathbb{E}[U]=\mu$. So,
$f_{\text{opt}}(X_{h})=AX_{h}+B\mu.$
The prediction error is $\epsilon=X_{f}-f(X_{h})$. Substituting $X_{f}$:
$\epsilon=(AX_{h}+BU)-f(X_{h}).$
We can rewrite this by adding and subtracting $B\mu$:
$\epsilon=\underbrace{(AX_{h}+B\mu-f(X_{h}))}_{\text{Model Inadequacy Term }\Delta_{f}(X_{h})}+\underbrace{B(U-\mu)}_{\text{Zero-Mean Stochastic Term}}.$ (22)
Let $\Delta_{f}(X_{h})=(AX_{h}+B\mu-f(X_{h}))$. This term represents how well the model $f(X_{h})$ approximates the optimal predictor $f_{\text{opt}}(X_{h})$. The stochastic term $B(U-\mu)$ has $\mathbb{E}[B(U-\mu)]=0$.
The mean of the error is $\mathbb{E}[\epsilon]=\mathbb{E}[\Delta_{f}(X_{h})]$. For an unbiased $f(X_{h})$ relative to $f_{opt}(X_{h})$, $\mathbb{E}[\Delta_{f}(X_{h})]=0$. The covariance of the error is $\text{Cov}(\epsilon)=\mathbb{E}[(\epsilon-\mathbb{E}[\epsilon])(\epsilon-\mathbb{E}[\epsilon])^{\top}]$. If we assume $\mathbb{E}[\Delta_{f}(X_{h})]=0$ (i.e., $f(X_{h})$ is unbiased for $AX_{h}+B\mu$ on average), then $\mathbb{E}[\epsilon]=0$. The error covariance is:
$\text{Cov}(\epsilon)=\mathbb{E}[\epsilon\epsilon^{\top}]=\mathbb{E}\left[(\Delta_{f}(X_{h})+B(U-\mu))(\Delta_{f}(X_{h})+B(U-\mu))^{\top}\right].$
Expanding this:
$\text{Cov}(\epsilon)=\mathbb{E}[\Delta_{f}(X_{h})\Delta_{f}(X_{h})^{\top}]+\mathbb{E}[\Delta_{f}(X_{h})(U-\mu)^{\top}B^{\top}]+\mathbb{E}[B(U-\mu)\Delta_{f}(X_{h})^{\top}]+\mathbb{E}[B(U-\mu)(U-\mu)^{\top}B^{\top}].$
Since $U\perp X_{h}$, $\Delta_{f}(X_{h})$ (a function of $X_{h}$) is independent of $U-\mu$. Thus, the cross-terms are zero:
$\mathbb{E}[\Delta_{f}(X_{h})(U-\mu)^{\top}B^{\top}]=\mathbb{E}[\Delta_{f}(X_{h})]\mathbb{E}[(U-\mu)^{\top}]B^{\top}=\mathbb{E}[\Delta_{f}(X_{h})]\cdot 0\cdot B^{\top}=0.$
So, the error covariance becomes:
$\text{Cov}(\epsilon)=\underbrace{\mathbb{E}[\Delta_{f}(X_{h})\Delta_{f}(X_{h})^{\top}]}_{\text{MSE of model inadequacy}}+B\Sigma B^{\top}.$ (23)
The term $\mathbb{E}[\Delta_{f}(X_{h})\Delta_{f}(X_{h})^{\top}]$ is the mean squared error of $f(X_{h})$ in approximating $AX_{h}+B\mu$. This term is always positive semi-definite. Therefore,
$\text{Cov}(\epsilon)\succeq B\Sigma B^{\top}.$
This means $B\Sigma B^{\top}$ is an irreducible lower bound on the error covariance, regardless of the complexity of $f(X_{h})$, as long as $f(X_{h})$ only uses $X_{h}$. If $f(X_{h})$ perfectly fits the optimal deterministic component, i.e., $f(X_{h})=f_{\text{opt}}(X_{h})=AX_{h}+B\mu$, then $\Delta_{f}(X_{h})=0$. The error reduces to:
$\epsilon=B(U-\mu).$ (24)
The covariance of this minimal error is:
$\text{Cov}(\epsilon)=B\Sigma B^{\top}.$ (25)
Any claim that a model $f^{\prime}(X_{h})$ achieves $\text{Cov}(\epsilon)\prec B\Sigma B^{\top}$ would imply that $\mathbb{E}[\Delta_{f^{\prime}}(X_{h})\Delta_{f^{\prime}}(X_{h})^{\top}]$ in Eq. (23) would have to be negative definite, which is impossible as it is a matrix of expected outer products (a sum of positive semi-definite matrices) which contradict with the independent intervention assumption. ∎
B.1.3 Real Scenario: Nonlinear Model, Nonlinear System
Proof.
Consider a general nonlinear system:
$X_{f}=F(X_{h},U),\quad U\sim\mathcal{P}_{U}\ \text{(i.i.d.)},\ \mathbb{E}[U]=\mu,\ \text{Cov}(U)=\Sigma,$ (26)
where $F$ is a nonlinear state transition function. We use a self-stimulated model:
$\hat{X}_{f}=f(X_{h}),$ (27)
where $f$ is an arbitrary nonlinear model.
The optimal model $f_{\text{opt}}(X_{h})$ that minimizes MSE is $\mathbb{E}[X_{f}\mid X_{h}]$. Assuming $U\perp X_{h}$:
$f_{\text{opt}}(X_{h})=\mathbb{E}_{U}[F(X_{h},U)\mid X_{h}]=\mathbb{E}_{U}[F(X_{h},U)].$
Let $F^{*}(X_{h})=\mathbb{E}_{U}[F(X_{h},U)]$. The prediction error is:
$\epsilon=X_{f}-f(X_{h})=\underbrace{(F^{*}(X_{h})-f(X_{h}))}_{\text{Model Inadequacy}}+\underbrace{(F(X_{h},U)-F^{*}(X_{h}))}_{\text{Irreducible Stochastic Error}}.$ (28)
The term $F(X_{h},U)-F^{*}(X_{h})$ has zero mean conditional on $X_{h}$ (and thus zero unconditional mean). The Model Inadequacy term $F^{*}(X_{h})-f(X_{h})$ reflects how well $f(X_{h})$ approximates the true conditional mean $F^{*}(X_{h})$.
The covariance of the error, assuming $\mathbb{E}[F^{*}(X_{h})-f(X_{h})]=0$, is:
$\text{Cov}(\epsilon)=\mathbb{E}[(F^{*}(X_{h})-f(X_{h}))(F^{*}(X_{h})-f(X_{h}))^{\top}]+\mathbb{E}[(F(X_{h},U)-F^{*}(X_{h}))(F(X_{h},U)-F^{*}(X_{h}))^{\top}].$
The cross-terms vanish due to the independence of $U$ from $X_{h}$ and the property that $\mathbb{E}_{U}[F(X_{h},U)-F^{*}(X_{h})\mid X_{h}]=0$. The first term is positive semi-definite. So,
$\text{Cov}(\epsilon)\succeq\mathbb{E}[(F(X_{h},U)-F^{*}(X_{h}))(F(X_{h},U)-F^{*}(X_{h}))^{\top}]=\mathbb{E}_{X_{h}}[\text{Cov}_{U}(F(X_{h},U)\mid X_{h})].$
The term $\text{Cov}_{U}(F(X_{h},U)\mid X_{h})$ is the conditional variance of $F(X_{h},U)$ given $X_{h}$. Using a first-order Taylor expansion for $F(X_{h},U)$ around $U=\mu$: $F(X_{h},U)\approx F(X_{h},\mu)+\nabla_{U}F(X_{h},\mu)(U-\mu).$ Then $F^{*}(X_{h})=\mathbb{E}_{U}[F(X_{h},U)]\approx F(X_{h},\mu)+\nabla_{U}F(X_{h},\mu)\mathbb{E}_{U}[U-\mu]=F(X_{h},\mu)$, neglecting higher-order terms (e.g., terms like $\frac{1}{2}\text{Tr}(\Sigma H_{F})$ where $H_{F}$ is the Hessian w.r.t $U$). Under this approximation, the irreducible stochastic error is $F(X_{h},U)-F^{*}(X_{h})\approx\nabla_{U}F(X_{h},\mu)(U-\mu)$. The conditional error covariance, given $X_{h}$, is approximately:
$\text{Cov}(\epsilon\mid X_{h};f=f_{\text{opt}})\approx\mathbb{E}_{U}[\nabla_{U}F(X_{h},\mu)(U-\mu)(U-\mu)^{\top}\nabla_{U}F(X_{h},\mu)^{\top}\mid X_{h}].$ (29)
Since $U\perp X_{h}$, this becomes $\nabla_{U}F(X_{h},\mu)\Sigma\nabla_{U}F(X_{h},\mu)^{\top}$. Higher-order terms in the Taylor expansion of $F$ would contribute terms of $\mathcal{O}(\Sigma^{2})$ etc.
For general nonlinear systems, the unconditional error covariance of the optimal model $f_{opt}(X_{h})$ satisfies (using this first-order approximation for the conditional covariance):
$\text{Cov}(\epsilon)\succeq\mathbb{E}_{X_{h}}\left[\nabla_{U}F(X_{h},\mu)\Sigma\nabla_{U}F(X_{h},\mu)^{\top}\right].$ (30)
This lower bound reflects the inherent system stochasticity due to $U$ and its propagation through the system dynamics $\nabla_{U}F$. ∎
B.1.4 Justification of Proposition B.1: Universality of the Self-Stimulation Error Floor
Proposition B.1 (Self-Stimulation Error Floor).
For any self-stimulated model $\hat{X}_{f}=f(X_{h})$ applied to a system $X_{f}=F(X_{h},U)$ where $U\perp X_{h}$, $\mathbb{E}[U]=\mu$, $\text{Cov}(U)=\Sigma$, the error covariance $\text{Cov}(\epsilon)=\mathbb{E}[(\epsilon-\mathbb{E}[\epsilon])(\epsilon-\mathbb{E}[\epsilon])^{\top}]$ (or $\mathbb{E}[\epsilon\epsilon^{\top}]$ if $\mathbb{E}[\epsilon]=0$) satisfies the following lower bound:
$\text{Cov}(\epsilon)\succeq\mathbb{E}_{X_{h}}\left[\text{Cov}_{U}(F(X_{h},U)\mid X_{h})\right].$ (31)
Using a first-order approximation $\text{Cov}_{U}(F(X_{h},U)\mid X_{h})\approx\nabla_{U}F(X_{h},\mu)\Sigma\nabla_{U}F(X_{h},\mu)^{\top}$, this becomes:
$\text{Cov}(\epsilon)\succeq\mathbb{E}_{X_{h}}\left[\nabla_{U}F(X_{h},\mu)\Sigma\nabla_{U}F(X_{h},\mu)^{\top}\right].$ (32)
Justification of Proposition B.1.
This proposition highlights that the self-stimulation error floor arises from fundamental system properties.
1. Intrinsic Limitation of Self-Stimulation: The error floor is fundamentally caused by the model’s inability to account for the specific realization of the stochastic intervention $U$, as it only has access to $X_{h}$. Even if the model $f(X_{h})$ perfectly learns the true conditional mean behavior of the system, i.e., $f(X_{h})=\mathbb{E}_{U}[F(X_{h},U)\mid X_{h}]$, the inherent variability of $F(X_{h},U)$ around this mean, $F(X_{h},U)-\mathbb{E}_{U}[F(X_{h},U)\mid X_{h}]$, introduces an irreducible noise component whose variance cannot be eliminated by any function of $X_{h}$ alone.
2. Generalization Across System Classes:
  • Linear Systems: If $F(X_{h},U)=AX_{h}+BU$, then $\mathbb{E}_{U}[F(X_{h},U)\mid X_{h}]=AX_{h}+B\mu$. The irreducible error term is $B(U-\mu)$. Its covariance is $B\Sigma B^{\top}$. The gradient $\nabla_{U}F(X_{h},\mu)=B$. The right-hand side of Eq. (32) becomes $\mathbb{E}_{X_{h}}[B\Sigma B^{\top}]=B\Sigma B^{\top}$, matching the exact result for linear systems.
  • Nonlinear Systems: If $F(X_{h},U)$ is nonlinear, the exact irreducible error covariance is $\mathbb{E}_{X_{h}}[\text{Cov}_{U}(F(X_{h},U)\mid X_{h})]$. The approximation $\mathbb{E}_{X_{h}}\left[\nabla_{U}F(X_{h},\mu)\Sigma\nabla_{U}F(X_{h},\mu)^{\top}\right]$ captures the first-order effect of $U$’s variance. The bound’s magnitude depends on the structure of $F$, but a lower bound due to $U$ always holds.
3. Independence-Driven Irreducibility: The independence $U\perp X_{h}$ is crucial. It implies that $\mathbb{E}_{U}[F(X_{h},U)\mid X_{h}]=\mathbb{E}_{U}[F(X_{h},U)]$, and it ensures that the error covariance decomposes additively. Let $f_{opt}(X_{h})=\mathbb{E}_{U}[F(X_{h},U)\mid X_{h}]$. The total error is $\epsilon=(F(X_{h},U)-f_{opt}(X_{h}))+(f_{opt}(X_{h})-f(X_{h}))$. The covariance $\text{Cov}(\epsilon)$ is the sum of the covariances of these two terms because the cross-term vanishes:
$\mathbb{E}\left[(f_{opt}(X_{h})-f(X_{h}))(F(X_{h},U)-f_{opt}(X_{h}))^{\top}\right]$
$=\mathbb{E}_{X_{h}}\left[(f_{opt}(X_{h})-f(X_{h}))\mathbb{E}_{U}[(F(X_{h},U)-f_{opt}(X_{h}))^{\top}\mid X_{h}]\right]=0,$
since $\mathbb{E}_{U}[F(X_{h},U)-f_{opt}(X_{h})\mid X_{h}]=0$ by definition of $f_{opt}$. The term $\text{Cov}(F(X_{h},U)-f_{opt}(X_{h}))$ is the irreducible part.
Implications of Proposition B.1: This provides a definitive justification for the existence of an error floor:
  • Self-stimulated models $f(X_{h})$ are fundamentally constrained to predicting the conditional expectation $\mathbb{E}[X_{f}\mid X_{h}]$. They cannot predict the specific deviation from this mean caused by the unobserved realization of $U$.
  • The error floor, characterized by $\mathbb{E}_{X_{h}}[\text{Cov}_{U}(F(X_{h},U)\mid X_{h})]$ (and approximated by Eq. (32)), is fundamental. It arises from the system’s inherent stochastic properties due to $U$ and its sensitivity to $U$, not from any particular choice of model $f(X_{h})$ (assuming $f(X_{h})$ can at best learn $\mathbb{E}[X_{f}|X_{h}]$).
  • To reduce or eliminate this error floor, it is necessary to gain information about $U$, for example, by incorporating external measurements related to $U$ or by explicitly modeling $U$’s dynamics if possible, thus going beyond simple self-stimulation based on $X_{h}$ alone.
B.2 Proposition 3.1: Intervention Efficacy
Proof.
Consider a system with $p$ independent interventions:
[…]
Let the true dynamics be $X_{f}=F(X_{h},U_{t})$, and let the model incorporate a subset of known interventions $\{U_{j}^{t}\}$:
$\hat{X}_{f}=f_{\theta}(X_{h},U_{j}^{t}).$ (34)
The prediction error is:
$\epsilon=X_{f}-\hat{X}_{f}=F(X_{h},U_{t})-f_{\theta}(X_{h},U_{j}^{t}).$ (35)
Now, decompose $U_{t}$ into known ($U_{j}^{t}$) and unknown ($U_{-j}^{t}$) components:
$\epsilon=\underbrace{F(X_{h},U_{j}^{t},U_{-j}^{t})-F(X_{h},U_{j}^{t},\mu_{-j})}_{\text{Reducible Error}}+\underbrace{F(X_{h},U_{j}^{t},\mu_{-j})-f_{\theta}(X_{h},U_{j}^{t})}_{\text{Model Mismatch}},$ (36)
where $\mu_{-j}=\mathbb{E}[U_{-j}^{t}]$.
Next, under optimal training, $f_{\theta}$ minimizes the mean squared error. This forces:
$f_{\theta}^{*}(X_{h},U_{j}^{t})=\mathbb{E}_{U_{-j}^{t}}[F(X_{h},U_{j}^{t},U_{-j}^{t})\mid X_{h},U_{j}^{t}].$ (37)
The reducible error then simplifies to:
$\epsilon=F(X_{h},U_{j}^{t},U_{-j}^{t})-F(X_{h},U_{j}^{t},\mu_{-j}).$ (38)
Now, consider the covariance reduction analysis:
1. Linear Systems: For $F(X_{h},U_{t})=AX_{h}+\sum_{i=1}^{p}B_{i}U_{i}^{t}$, the prediction error becomes:
$\epsilon=\sum_{i\neq j}B_{i}(U_{i}^{t}-\mu_{i}).$ (39)
The error covariance reduces by:
$\Delta\text{Cov}(\epsilon)=B_{j}\Sigma_{j}B_{j}^{\top}.$ (40)
2. Nonlinear Systems: For general $F(X_{h},U_{t})$, approximate via Taylor expansion at $U_{-j}^{t}=\mu_{-j}$:
$\epsilon\approx\nabla_{U_{-j}}F(X_{h},U_{j}^{t},\mu_{-j})(U_{-j}^{t}-\mu_{-j}).$ (41)
The covariance reduction becomes:
$\Delta\text{Cov}(\epsilon)=\nabla_{U_{j}}F\Sigma_{j}(\nabla_{U_{j}}F)^{\top}.$ (42)
Next, the independence argument:
The independence $U_{j}^{t}\perp U_{-j}^{t}$ ensures:
$\text{Cov}(\epsilon)=\text{Cov}(\text{Reducible Error})+\text{Cov}(\text{ModelMismatch}).$ (43)
Optimal training nullifies the model mismatch term, leaving:
$\text{Cov}(\epsilon)\succeq\sum_{i\neq j}\nabla_{U_{i}}F\Sigma_{i}(\nabla_{U_{i}}F)^{\top}.$ (44)
This matches Proposition 3.1’s claim.
Finally, we align with Proposition 2.1. The irreducible error floor in Proposition 2.1 is partially "carved out" by incorporating $U_{j}^{t}$. The reduction $\Delta\text{Cov}(\epsilon)$ quantifies how much intervention knowledge lifts the theoretical performance ceiling.
This concludes the justification that Proposition 3.1 rigorously formalizes the intuition that any measurable intervention knowledge reduces forecasting uncertainty, even under partial observability. ∎
B.2.1 Case Study: Dual-Intervention Linear System
System Setup
Consider a linear system with two independent interventions:
$X_{f}=AX_{h}+B_{1}U_{1}+B_{2}U_{2},\quad U_{1}\sim\mathcal{N}(0,\sigma_{1}^{2}),\ U_{2}\sim\mathcal{N}(0,\sigma_{2}^{2}),$ (45)
where:
$A=\begin{bmatrix}0.8&0\\ 0&0.8\end{bmatrix},\ B_{1}=\begin{bmatrix}1\\ 0\end{bmatrix},\ B_{2}=\begin{bmatrix}0\\ 1\end{bmatrix},\ \sigma_{1}^{2}=0.5,\ \sigma_{2}^{2}=0.3.$ (46)
The self-stimulated baseline model is:
$\hat{X}_{f}^{\text{(base)}}=CX_{h}+d.$ (47)
Case 1: No Intervention Knowledge
Using least squares, the optimal parameters are:
$C^{*}=A,\quad d^{*}=B_{1}\mu_{1}+B_{2}\mu_{2}=0\quad(\text{since }\mu_{1}=\mu_{2}=0).$ (48)
Prediction error:
$\epsilon^{\text{(base)}}=B_{1}U_{1}+B_{2}U_{2}.$ (49)
Error covariance:
$\text{Cov}(\epsilon^{\text{(base)}})=B_{1}\sigma_{1}^{2}B_{1}^{\top}+B_{2}\sigma_{2}^{2}B_{2}^{\top}=\begin{bmatrix}0.5&0\\ 0&0.3\end{bmatrix}.$ (50)
Case 2: Partial Intervention Knowledge (Observing $U_{1}$)
Extend the model to leverage $U_{1}$:
$\hat{X}_{f}^{\text{(IATSF)}}=AX_{h}+B_{1}U_{1}+d.$ (51)
Optimal bias term:
$d^{*}=B_{2}\mu_{2}=0.$ (52)
Prediction error:
$\epsilon^{\text{(IATSF)}}=B_{2}(U_{2}-\mu_{2})=B_{2}U_{2}.$ (53)
Error covariance:
$\text{Cov}(\epsilon^{\text{(IATSF)}})=B_{2}\sigma_{2}^{2}B_{2}^{\top}=\begin{bmatrix}0&0\\ 0&0.3\end{bmatrix}.$ (54)
The error is reduced by:
$\Delta\text{Cov}(\epsilon)=\begin{bmatrix}0.5&0\\ 0&0\end{bmatrix}=B_{1}\sigma_{1}^{2}B_{1}^{\top}.$ (55)
This matches Proposition 3.1 for linear systems.
Case3: Nonlinear Extension
For a weakly nonlinear system $X_{f}=AX_{h}+\sin(U_{1})B_{1}+U_{2}B_{2}$:
  • Unknown $U_{1},U_{2}$:
    $\text{Cov}(\epsilon)\approx B_{1}\cos^{2}(\mu_{1})\sigma_{1}^{2}B_{1}^{\top}+B_{2}\sigma_{2}^{2}B_{2}^{\top}=\begin{bmatrix}0.5\cos^{2}(0)&0\\ 0&0.3\end{bmatrix}.$ (56)
  • Known $U_{1}$:
    $\text{Cov}(\epsilon)\approx B_{2}\sigma_{2}^{2}B_{2}^{\top}=\begin{bmatrix}0&0\\ 0&0.3\end{bmatrix}.$ (57)
Conclusion: Even in nonlinear systems, measurable interventions reduce the error bound by their sensitivity-weighted variance, as formalized in Proposition 3.1.
B.3 Error Introduced by Intervention Forecaster and Benchmark Design
B.3.1 Error Propagation with Non-Optimizable Intervention Forecasting
Consider a linear system with historical state $X_{h}$ and future intervention $U_{f}$:
[…]
where $U_{f}$ impacts the system instantaneously. Thus, - In training Phase: Uses true historical-future pairs $(X_{h},X_{f},U_{f})$. - Testing Phase: Requires forecasting $U_{f}$ externally. The forecaster is fixed (not optimizable) and produces:
$\hat{U}_{f}=U_{f}+\epsilon_{f},\quad\epsilon_{f}\sim\mathcal{N}(0,\Sigma_{\hat{U}}).$ (59)
After training, the model $\hat{X}_{f}=AX_{h}+BU_{f}$ achieves zero error if $\Sigma_{w}=0$:
$\epsilon_{\text{train}}=X_{f}-\hat{X}_{f}=w_{h}\quad\Rightarrow\quad\text{Cov}(\epsilon_{\text{train}})=\Sigma_{w}.$ (60)
In testing, predictions use the fixed forecaster $\hat{U}_{f}$:
$\hat{X}_{f}=AX_{h}+B\hat{U}_{f}=AX_{h}+B(U_{f}+\epsilon_{f}).$ (61)
The prediction error becomes:
$\epsilon_{\text{test}}=X_{f}-\hat{X}_{f}=\underbrace{w_{h}}_{\text{System Noise}}-\underbrace{B\epsilon_{f}}_{\text{Irreducible Forecaster Error}}.$ (62)
Error covariance (assuming $w_{h}\perp\epsilon_{f}$):
$\text{Cov}(\epsilon_{\text{test}})=\Sigma_{w}+B\Sigma_{\hat{U}}B^{\top}.$ (63)
Thus, we find that, 1. The term $B\Sigma_{\hat{U}}B^{\top}$ dominates if $\Sigma_{w}\ll B\Sigma_{\hat{U}}B^{\top}$. This error is independent of model quality. 2. Since $\Sigma_{\hat{U}}$ is fixed and external, test error does not reflect the model’s inherent capability. A "good" model may appear poor due to a low-quality forecaster.
Case Study: Separating Model and Forecaster Effects
Let $A=I$, $B=I$, $\Sigma_{w}=0$, and $\Sigma_{\hat{U}}=0.5I$, we can train a perfect model: $\hat{X}_{f}=X_{h}+\hat{U}_{f}$.
But its test error gives:
$\text{Cov}(\epsilon_{\text{test}})=0+I\cdot 0.5I\cdot I^{\top}=0.5I.$ (64)
Despite a perfect model, test error is entirely dictated by $\Sigma_{\hat{U}}$.
Final Conclusion: When interventions are forecasted by a non-optimizable external module, the test error upper bound is fundamentally constrained by:
$\text{Cov}(\epsilon_{\text{test}})\succeq B\Sigma_{\hat{U}}B^{\top}$ (65)
This invalidates isolated model evaluation—performance metrics inherently conflate model and forecaster limitations.
B.3.2 Perfect Intervention Forecaster Assumption for Fairness
To eradicate the noise introduced by the inaccurate intervention forecaster for a fair benchmarking we assume that we have a perfect intervention forecaster.
Assumption of Accurate Forecaster
Assume the intervention forecaster is highly accurate, with negligible error:
$\Sigma_{\hat{U}}\approx 0\quad\Rightarrow\quad\hat{U}_{f}\approx U_{f}.$ (66)
In this idealized scenario, the test-time prediction error reduces to:
$\text{Cov}(\epsilon_{\text{test}})=\Sigma_{w}+B\cdot 0\cdot B^{\top}=\Sigma_{w}.$ (67)
Implications for Model Evaluation
1. Fair Assessment: With $\Sigma_{\hat{U}}\approx 0$, the test error $\text{Cov}(\epsilon_{\text{test}})=\Sigma_{w}$ directly reflects the model’s inherent capability, as it matches the training error bound.
2. Decoupling Forecaster Effects: A perfect forecaster eliminates the confounding term $B\Sigma_{\hat{U}}B^{\top}$, isolating the model’s performance. This allows direct comparison between different models or training methodologies.
3. Revealing True Limitations: Any residual error $\Sigma_{w}$ now purely represents: - Fundamental system noise (unavoidable), - Model limitations (e.g., parameter estimation errors, structural mismatch).
B.4 Error Introduced by Weight Sharing
B.4.1 Linear System Analysis
Consider a linear observation model with historical state $Z_{h}$ and multi-channel observations:
$X_{f}=CZ_{h}=\begin{bmatrix}C_{1}Z_{h}\\ C_{2}Z_{h}\\ \vdots\\ C_{k}Z_{h}\end{bmatrix},\quad C_{i}\in\mathbb{R}^{1\times n},$ (68)
where $C_{i}$ is the distinct observation matrix for channel $i$.
Assume all channels share a single weight $c\in\mathbb{R}^{1\times n}$:
$\hat{X}_{f}=\mathbf{1}_{k}\cdot cZ_{h}=\begin{bmatrix}cZ_{h}\\ cZ_{h}\\ \vdots\\ cZ_{h}\end{bmatrix}.$ (69)
The prediction error becomes:
$\epsilon=X_{f}-\hat{X}_{f}=\begin{bmatrix}(C_{1}-c)Z_{h}\\ (C_{2}-c)Z_{h}\\ \vdots\\ (C_{k}-c)Z_{h}\end{bmatrix}=(C-\mathbf{1}_{k}c)Z_{h}.$ (70)
Let $\Sigma_{Z}=\text{Cov}(Z_{h})$. The error covariance is:
$\text{Cov}(\epsilon)=(C-\mathbf{1}_{k}c)\Sigma_{Z}(C-\mathbf{1}_{k}c)^{\top}.$ (71)
The optimal shared weight $c_{\text{opt}}$ minimizing the trace is:
$c_{\text{opt}}=\frac{1}{k}\sum_{i=1}^{k}C_{i}.$ (72)
Substituting $c_{\text{opt}}$, the irreducible error covariance becomes:
$\text{Cov}(\epsilon_{\text{opt}})=\left(C-\frac{1}{k}\mathbf{1}_{k}\mathbf{1}_{k}^{\top}C\right)\Sigma_{Z}\left(C-\frac{1}{k}\mathbf{1}_{k}\mathbf{1}_{k}^{\top}C\right)^{\top}.$ (73)
B.4.2 Nonlinear System Generalization
For nonlinear observations $X_{f}=\mathcal{O}(Z_{h})=[o_{1}(Z_{h}),\dots,o_{k}(Z_{h})]^{\top}$, a weight-shared model forces:
$\hat{X}_{f}=\mathbf{1}_{k}\cdot o(Z_{h}).$ (74)
The error is:
$\epsilon=\begin{bmatrix}o_{1}(Z_{h})-o(Z_{h})\\ \vdots\\ o_{k}(Z_{h})-o(Z_{h})\end{bmatrix}.$ (75)
Assume $o_{i}(Z_{h})=o(Z_{h})+\Delta_{i}(Z_{h})$ with $\Delta_{i}\sim\mathcal{N}(0,\Sigma_{i})$. The covariance becomes:
$\text{Cov}(\epsilon)=\text{diag}(\Sigma_{1},\dots,\Sigma_{k}).$ (76)
B.4.3 Justification: Key Analogies to Previous Framework
  • Structural Bias: Weight-sharing corresponds to assuming $U_{f}$ is constant across channels, analogous to ignoring external interventions.
  • Irreducible Error: The term $\text{Cov}(\epsilon_{\text{opt}})$ mirrors $B\Sigma B^{\top}$, where $\Sigma$ represents unmodeled channel-specific variations.
  • Sensitivity Amplification: The matrix $C-\mathbf{1}_{k}c_{\text{opt}}$ amplifies discrepancies, similar to $\nabla_{U}F$ in nonlinear systems.
B.4.4 Case Study: Two Channels
Let $o_{1}(Z_{h})=Z_{h}$, $o_{2}(Z_{h})=2Z_{h}$, and force $o(Z_{h})=aZ_{h}$. The optimal $a=1.5$ yields:
$\text{Cov}(\epsilon)=\frac{1}{4}\text{Var}(Z_{h})\begin{bmatrix}1&1\\ 1&1\end{bmatrix}.$ (77)
Conclusion: Weight-sharing introduces an error floor governed by:
$\text{Cov}(\epsilon)\succeq\left(C-\frac{1}{k}\mathbf{1}_{k}\mathbf{1}_{k}^{\top}C\right)\Sigma_{Z}\left(C-\frac{1}{k}\mathbf{1}_{k}\mathbf{1}_{k}^{\top}C\right)^{\top}$ (78)
This matches the structure of Proposition 2.1, where unmodeled channel diversity plays the role of unobserved interventions. Breaking this bound requires abandoning weight-sharing or introducing channel-specific adapters.
B.5 Error Introduced by Incomplete Observation
Finally, we would like to discuss the error introduced by incomplete observation which is also an inherent error source in the TSF. This shows that our given lower bound is already a very ideal and conservative, there is still a lot loophole in the TSF task formulation.
Consider a hidden state system with partial observations:
$Z_{f}=AZ_{h}+BU,\quad U\sim\mathcal{N}(\mu,\Sigma),$ (79)
where $Z_{h}\in\mathbb{R}^{n}$ is the historical hidden state. The observable state is:
$X_{h}=HZ_{h},\quad H\in\mathbb{R}^{m\times n},\ \text{rank}(H)=m<n.$ (80)
The observable dynamics become:
$X_{f}=HZ_{f}=HAZ_{h}+HBU.$ (81)
The hidden state can be decomposed as:
$Z_{h}=H^{+}X_{h}+\tilde{Z}_{h},$ (82)
where $H^{+}$ is the pseudo-inverse of $H$, and $\tilde{Z}_{h}$ represents the unobservable state component. Substituting into $X_{f}$:
$X_{f}=HAH^{+}X_{h}+HA\tilde{Z}_{h}+HBU.$ (83)
A self-stimulated model predicts:
$\hat{X}_{f}=CX_{h}+d.$ (84)
The prediction error is:
$\epsilon=X_{f}-\hat{X}_{f}=(HAH^{+}-C)X_{h}+HA\tilde{Z}_{h}+HB(U-\mu).$ (85)
The least squares solution gives:
$C^{*}=HAH^{+},\quad d^{*}=HB\mu.$ (86)
The irreducible error becomes:
$\epsilon=HA\tilde{Z}_{h}+HB(U-\mu).$ (87)
The error covariance splits into two components:
$\text{Cov}(\epsilon)=\underbrace{HA\text{Cov}(\tilde{Z}_{h})(HA)^{\top}}_{\text{Hidden State Error}}+\underbrace{HB\Sigma B^{\top}H^{\top}}_{\text{Intervention Error}}.$ (88)
Finally, the error lower bound comes from: 1. Hidden State Error: Propagates through $HA$ from the unobservable subspace, governed by $\text{Cov}(\tilde{Z}_{h})$. 2. Intervention Error: Matches Proposition 2.1’s bound $B\Sigma B^{\top}$, projected onto the observable space via $H$.
The total error lower bound becomes:
$\text{Cov}(\epsilon)\succeq HB\Sigma B^{\top}H^{\top}+HA\text{Cov}(\tilde{Z}_{h})(HA)^{\top}$ (89)
This extends Proposition 2.1 by adding a term from partial observability. The bound is conservative because:
  • Hidden state error $\text{Cov}(\tilde{Z}_{h})$ depends on system stability in the unobservable subspace.
  • Noisy interventions $U$ remain irreducible without direct measurement.
Appendix C Data and Code Availability
The code for FIATS is available at: https://anonymous.4open.science/r/IATSF_review-F624. Along with script for creating the toy and electricity dataset!
However, the Atmospheric Physics dataset is to large for anonymous sharing. We upload a sample for inspection.
We will finally release all the time series, raw text and pre-embedded text embedding after the anonymous review period.
Appendix D Related Works
D.1 Text Embedding Model
Text embedding models have undergone significant advancements, providing efficient and semantically rich vector representations of textual information. Early transformer-based models like BERT [25] encode sentences into embeddings by pretraining on masked language modeling tasks, enabling them to capture contextual semantics. However, BERT embeddings are not specifically optimized for tasks requiring fine-grained semantic similarity, prompting the development of more task-specific models.
MPNet [23] and MiniLM [22] build upon BERT  by introducing novel architectural and pretraining strategies. MPNet combines masked language modeling with permuted sequence prediction, allowing for better contextual understanding and token dependencies. MiniLM, on the other hand, employs knowledge distillation to create smaller, faster models that retain high performance, making them ideal for resource-constrained applications.
OpenAI’s embedding models [26] represent another major step forward, leveraging large-scale proprietary transformer architectures. These embeddings are designed to excel in tasks like semantic search, classification, and similarity, offering generalizability and strong performance across a variety of applications. They also incorporate dimensional flexibility, allowing embeddings to be truncated or adjusted based on application needs, as seen with the Matryoshka embedding technique. This technique allows embeddings to maintain their semantic integrity even when their dimensions are reduced, offering scalability and adaptability.
A key property of text embeddings is their compatibility with similarity measures like cosine similarity. By projecting text into a shared semantic space, cosine similarity enables the computation of semantic closeness between embeddings, making it a foundational operation for tasks like clustering, retrieval, and alignment between modalities. This capability is crucial in applications requiring robust generalization across diverse textual expressions.
Together, these advancements have expanded the utility of text embeddings in various domains, including information retrieval, natural language understanding, and multimodal learning tasks. Our work builds on these innovations by leveraging pre-trained text embeddings for aligning textual semantics with time series patterns, ensuring robust causal modeling and efficient text-guided time series forecasting.
D.2 Time Series Analysis with Text Embedding
Adding more information to time series by incorporating heterogeneous information has been a long-studied topic, with several works opting to use text embeddings as input.
In the financial field, where time series are often more correlated to external information, several works [27, 28] have used text embeddings as external graph relationships to capture the correlations between keywords and stock descriptions, further influencing the ranking process in stock trading. More recently, a line of works [15, 29] has sought to enrich time series data by adding news text embeddings to the time series embeddings. However, these methods still face limitations in solving information insufficiency, as they do not incorporate causal information that could guide the model in predicting time series patterns driven by external events. Additionally, these works primarily use external text embeddings to expand the lookback window, without fully exploiting the underlying properties of the text embeddings.
To tackle these challenges, we introduce the Time-Series Guided Text Forecasting (IATSF) model, which expands traditional time series forecasting by incorporating external textual data that offers causal insights. Unlike previous approaches that use text embeddings simply as supplementary information, IATSF leverages the text to provide causal guidance, aligning textual data with time series patterns. Through the integration of CASM, we can effectively extract channel-dynamic news correlations from the pre-trained text embeddings, enabling the model to adapt to the specific distributions of different time series channels. This allows the model to make more accurate predictions by incorporating both the semantic meaning of the text and its causal relationship with the time series data.
Appendix E About Predictability of Trend
In our study, we define "trend" as patterns that exhibit very low frequency while lacking periodicity within the observed time window, rather than simple exponential or linear patterns. For instance, the pressure channel in our Atmospheric Physics dataset exemplifies this with its irregular low-frequency fluctuations, which appear to be random and non-periodic. Such randomness hampers the model’s ability to learn stable patterns when relying solely on historical time series data.
However, these low-frequency patterns often correlate with external influences—for example, a drop in temperature due to cold air can significantly increase atmospheric pressure. By integrating this type of external information, our model is designed to discern causal relationships between such environmental factors and the observed low-frequency trends, thereby enhancing predictability.
For a practical illustration, please refer to the pressure (p-bar) channel in Fig. 4 and Fig. 7. This channel displays non-periodic fluctuations, which the traditional patchTST model even struggles produce a valid forecasting. In contrast, our IATSF model, which incorporates external textual cues, successfully tracks these changes, demonstrating the effectiveness of including external information for predicting complex trends.
Figure 7: Visualization on two channels with different time scale. Temperature channel shows obvious periodicity both daily and annually. However the atmosphere pressure seems to be noise on large time scale but shows slowly random changing low-frequency "trend". Which makes it hard to be predicted without external information.
Appendix F Experiment Settings
Toy Dataset
For Toy dataset, All of the models are following the same experimental setup with prediction length $H\in\{14,28,60,120\}$ and LBW length $T=60$.
Electrical Utility
For Electrical Utility dataset we follow the experiment settings in previous works as follows: forecasting horizon $H\in\{96,192,336,720\}$, look back window length of 288. For fair comparison, we directly compare with the results report in the baseline original paper. And the FIATS is trained with the captioned version. On this dataset, we also test the impact of a shorter look-back window on the FIATS.
Atmospheric Physic
We follow the experiment settings in previous works as follows: forecasting horizon $H\in\{96,192,336,720\}$, look back window length of 360. We trained all other models on the Atmospheric Physic 14-19 and 14-24 dataset with their setting on the original weather dataset accordingly. And the FIATS is trained with the captioned version.
GAUD
We split the dataset on each time series by 7:1:2 for training, validating and testing. We set the forecasting horizon H for 14 and look back window length of 60. For pretraining, we concatenate all the training set together to train and validate the model and test on each test set separately.
Appendix G Channel-wise performance on Atmospheric Physics Dataset
The difficulty in predicting each channel varies, therefore, we present channel-wise performance in Table 5. The results demonstrate that FIATS, with the aid of external textual climate reports, significantly enhances forecasting accuracy across all channels. Notably, the model achieves over a 60% performance improvement in channels such as atmospheric pressure (p (mbar)), relative humidity (rh(%)), and vapor pressure deficit (VPdef (mbar)), which typically cannot be predicted reliably using historical time series data alone. The integration of external text cues has led to groundbreaking improvements in forecasting these parameters.
However, the wind velocity channel shows minimal variance in performance across all models, each achieving similar results with slight losses. This phenomenon is attributed to the presence of extreme values in this channel, which, after normalization, diminish the impact of more typical values on the overall gradient. Consequently, all models struggle to learn detailed patterns in this channel due to the reduced contribution to the global gradient.
Another noteworthy observation is that while FIATS is capable of predicting rainfall—unlike models that default to predicting near-zero averages—the performance improvement in these channels is modest. This is because rainfall is relatively scarce in this dataset, leading to large losses when rain is inaccurately predicted at the wrong times. Conversely, predicting the average value results in a smaller overall loss. This tendency explains why other models often opt for the average, avoiding the complex task of learning rainfall patterns. Nevertheless, accurate rainfall forecasting remains crucial in meteorological applications, underscoring our commitment to enhancing predictive accuracy in this area. The same principle also applies to other channels.
Table 5: Channel wise performance on Weather-medium dataset in MSE. The best is highlighted in bold and the second best is highlighted in underline.
Channel FIATS FITS DLinear PatchTST iTransformer IMP.
p (mbar) 0.1365 0.8637 0.8238 0.9301 1.0320 83.43%
T (degC) 0.1889 0.2924 0.3329 0.2964 0.3233 35.40%
Tpot (K) 0.1829 0.3163 0.3525 0.3225 0.3533 42.18%
Tdew (degC) 0.3467 0.4043 0.4085 0.4082 0.4258 14.25%
rh (%) 0.2479 0.6541 0.6788 0.6997 0.8185 62.10%
VPmax (mbar) 0.2369 0.3500 0.3984 0.3521 0.4086 32.31%
VPact (mbar) 0.2998 0.3404 0.3534 0.3515 0.3845 11.93%
VPdef (mbar) 0.2835 0.6384 0.6968 0.6744 0.8038 55.59%
sh (g/kg) 0.2995 0.3434 0.3562 0.3557 0.3896 12.78%
H2OC (mmol/mol) 0.2996 0.3434 0.3562 0.3556 0.3894 12.75%
rho (g/m³) 0.1926 0.3909 0.4119 0.4182 0.4535 50.73%
wv (m/s) 0.0002 0.0002 0.0002 0.0002 0.0003 0.00%
max. wv (m/s) 0.0004 0.0005 0.0004 0.0005 0.0006 0.00%
wd (deg) 0.7270 1.1605 1.1295 1.1344 1.2735 35.64%
rain (mm) 0.6824 0.6905 0.7167 0.6891 0.6998 0.97%
raining (s) 0.7900 0.8735 0.9379 0.8591 0.9942 8.04%
SWDR (W/m²) 0.1828 0.3084 0.3856 0.2967 0.3776 38.39%
PAR (umol/m²/s) 0.1773 0.2840 0.3588 0.2704 0.3473 34.43%
max. PAR (umol/m²/s) 0.1975 0.2599 0.3195 0.2632 0.3226 24.01%
Tlog (degC) 0.1774 0.2802 0.3260 0.2806 0.3290 36.69%
CO2 (ppm) 0.2600 0.2716 0.2812 0.2618 0.2760 0.69%
Avg. Loss 0.2814 0.4317 0.4583 0.4391 0.4954 34.82%
Appendix H Performance Visualization on Atmospheric Physics Dataset
We provide the full visualization as Fig. 8. The FIATS shows great performance across all the channels. Even very hard ones such as Wind dir. It can also model the time series that totally independent with the weather such as the CO2 channel.
Figure 8: Full visualization all channels on the 15000th test sample of Weather-Caption-Medium dataset. Blue line for ground truth,Red line for FIATS, Green line for PatchTST and Orange line for FIATS with swapping the news on the second and forth forecasting day.
Appendix I Weather Results w. w/o. RIN
We compared the performance of models with and without Reversible Instance Normalization (RIN) on the Atmospheric Physics-medium dataset, focusing on a 720-hour forecasting horizon. The model with RIN enabled achieved an MSE of 0.3428, whereas the model without RIN achieved a lower MSE of 0.2814. Results visualized in Fig. 9 show that the RIN-enabled model exhibits significant biases in many channels, particularly those with gradual trend shifts. This occurs because RIN removes the bias term from all instances, leaving the model unable to recognize relative bias and trend values. For instance, with RIN, temperature patterns in winter and summer are treated similarly, ignoring the typically higher and more variable temperatures in summer. Additionally, we noted pronounced shifting behavior coinciding with changes in captions, suggesting that the absence of bias information leads the model to over-rely on textual prompts, compensating for the missing data.
Figure 9: Full visualization all channels on the 15000th test sample of Weather-Caption-Medium dataset. Blue line for ground truth, Red line for FIATS without RIN, Green line for FIATS with RIN enabled.
Appendix J Ablation Study on Causal Relationship Extraction
FIATS is designed to learn causal relationships between events described in text and their corresponding time series patterns. While not explicitly an alignment model, it effectively aligns the semantic meaning of text with the time series data it impacts. The model generates time series patterns guided by the textual information, and its performance varies based on the quality of the text input:
  1. 1.
    Training with Meaningful and Relevant Text:
    • Inference with Similar Text: Produces strong results by accurately extracting causal relationships between events in the text and time series patterns.
  2. 2.
    Training with Zero/Random Text:
    • Inference with Any Text: Produces results equivalent to PatchTST, as no additional information is present in the text. The model relies solely on the time series data, ignoring the random text.
  3. 3.
    Training with Meaningful Text, Inference with Incorrect Text:
    • Inference with Incorrect Text: Results are poor, as the model relies on the misleading text input and generates patterns based on incorrect or irrelevant information.
We detail the FIATS performance under different text conditions in the following table:
Table 6: FIATS performance under different training and testing conditions. We report the result of forecasting horizon 96 on Atmospheric Physics-medium dataset using MiniLM embedding.
Train with
Good Zero Random
Test with Good
0.186
(captures causal relationships)
0.249
(corrupted random patterns)
0.251
(similar to PatchTST)
Zero
0.724
(corrupted repetive patterns)
0.249
(similar to PatchTST)
0.254
(similar to PatchTST)
Random
0.615
(corrupted random patterns)
0.249
(similar to PatchTST)
0.250
(similar to PatchTST)
The results of the ablation study provide strong evidence that FIATS relies on capturing causal relationships between time series patterns and dynamic news, rather than simply treating text as auxiliary input. When trained with meaningful and correlated news, the model demonstrates its ability to effectively extract these relationships, yielding strong predictive performance. This highlights FIATS’s capacity to align the semantic meaning of text with time series patterns in a causally meaningful way.
On the other hand, when trained with good text but tested with random or misleading text, the model produces poor predictions because it continues to rely on the input text, even when it is inaccurate or irrelevant. This further underscores the model’s dependence on the quality of the textual input rather than merely defaulting to learned time series patterns.
Interestingly, when trained with bad or random text, FIATS fails to establish causal relationships and instead reverts to PatchTST-level performance, indicating it falls back to relying solely on time series data. Furthermore, when subsequently tested with good text, the model trained on bad text still ignores the input entirely, suggesting it stops depending on textual input when the training data lacks meaningful causal relationships.
These results collectively demonstrate that FIATS’s strength lies in its ability to extract and leverage causal relationships between text and time series data. The model’s performance is tightly coupled with the quality and relevance of the textual input, validating the centrality of causal alignment in its design and functionality.
These outcomes demonstrate that IATSF effectively achieves alignment in the “event” space, linking events described in the text to the corresponding time series patterns.
Appendix K Attention Map Visualization on Atmospheric Physics
We further visualize two cross-attention blocks to further investigate the FIATS. You are strongly advised to check the Tab. 13, Appendix P.4.5 and Fig. 8 while reading this part.
Figure 10 illustrates the attention map of the "text-guided channel independent" cross-attention block in the text encoder across three layers. In the first layer, attention is predominantly focused on the first sentence, which specifies the month and time. This sentence is crucial as it provides temporal context that significantly impacts the prediction of both daily and annual periodicity. While other sentences receive moderate attention, the sixth sentence, which describes atmospheric pressure as detailed in Table 13, consistently receives no attention across all channels.
In the second layer, however, there is a notable shift in attention dynamics. All channels, particularly channel 0, show intense focus on the sixth sentence. According to the channel definitions in Appendix P.4.5, channel 0 directly corresponds to atmospheric pressure. Channels 10 and 20, which are related to air density and CO2 concentration respectively—factors closely associated with pressure—also display relatively high attention scores. This suggests that the FIATS is capable of discerning the underlying relationships among the channels.
The separation of attention focus between the first and second layers suggests that the influence of atmospheric pressure on the model’s predictions is independent of time. In the third layer, a diversity of attention patterns emerges; channel 0 focuses exclusively on the sixth sentence, while other channels predominantly attend to the first sentence.
Since we take the output of previous layer as query and input news embeddings as key and value, the information lies in the news are progressively added to the channel embeddings. Thus, the model can focus on different perspective in separate cross attention layers.
(a) CASM Layer1
(b) CASM Layer2
(c) CASM Layer3
Figure 10: Attention map of the "CASM" cross attention block on the 15000th test sample of Atmospheric Physics dataset dataset. We use three cross attention block. The vertical axis stand for channels and horizon stand for the 7 sentences of the weather report summary.
Figure 11 presents the attention map of the modality mixer layer cross attention block in the Atmospheric Physics dataset. The map, averaged across three cross attention layers, illustrates distinct patterns of attention for each channel. This diversity underscores the FIATS’s ability to adaptively extract time series embeddings tailored to the unique distribution characteristics of each channel, facilitated by textual inputs.
Notably, the channels for SWDR, PAR, and max.PAR display clear periodic patterns in their attention maps, aligning with observations from waveform visualizations. These patterns suggest that the FIATS effectively captures and utilizes periodic information from these environmental variables.
Furthermore, the channels labeled rain and raining show a particularly interesting behavior; they assign significantly higher attention scores to the exact time periods of rainfall within the look-back window. This behavior indicates that the FIATS is adept at identifying and prioritizing crucial temporal events specific to each channel, further enhancing its forecasting accuracy by focusing on relevant patterns where needed. This level of detail in attention allocation demonstrates the model’s capability to integrate contextual cues from textual data and further guide the time series forecasting.
Figure 11: Attention map of the modality mixer layer cross attention block on Atmospheric Physics dataset, on the 15000th test sample of Weather-Caption-Medium dataset. The attention map is averaged across three cross attention layers. We plot the attention map for each channel. The vertical axis stand for output time series patches and the horizon stand for the input time series patches embedding from PatchTST backbone.
Appendix L Comparison with More Baselines on Electrical Utility
We further compare with more baselines on Electrical Utility, including Autoformer, Fedformer, Informer, FiLM and TimesNet [30, 31, 16, 32, 33]. FIATS shows dominant superior performance across these baselines, as shown in Tab. 7.
Table 7: The comparison on Electricity dataset with other baselines. Best is marked in bold and the second best is marked in underline.
Pred. Len. FIATS FIATS_120 Autoformer Fedformer Informer FiLM TimesNet
96 0.124 0.127 0.201 0.188 0.274 0.154 0.168
192 0.144 0.146 0.222 0.197 0.296 0.164 0.184
336 0.16 0.164 0.231 0.212 0.3 0.188 0.198
Elec 720 0.193 0.200 0.254 0.244 0.373 0.236 0.22
Appendix M Full Result on GAUD Dataset
We show the comparison on GAUD Dataset in Tab. 8, and Tab, 9 with PatchTST. We use de-normed MAE as metric since the base volume of players of each game varies drastically, using normed metrics can lead to unfair comparison. The pretrain indicate that the model is jointly trained on all the games and each game is labeled by the channel discription. The Gnorm indicate that we apply the global normalization to preserve the player variation mentioned before. But it seems bring limited boost.
Table 8: Full result on GAUD dataset in de-normalized MAE. The best result is shown in green shaded bold font. The ones with performance boost over 10% is marked in red.
game_id IATSF_pretrain
IATSF_pretrain
_Gnorm
IATSF PatchTST IMP/%
10 781.2257 724.5196 849.8968506 804.757019 0.099704
240 372.887 374.03445 400.0899353 421.2349243 0.114777
440 10216.276 10816.979 10694.47852 10828.37891 0.056528
550 4110.673 4437.846 4336.002441 4587.186035 0.103879
570 30179.111 30914.955 31916.02344 32031.38477 0.057827
620 674.1199 789.6992 799.2894897 710.8320313 0.051647
730 51687.87 50937.348 52638.79297 51069.08984 0.00258
3590 687.4657 775.8215 734.3128662 743.4817505 0.075343
39210 3310.422 3719.8003 3388.064453 3420.974609 0.032316
105600 4415.245 4899.586 4464.801758 4513.034668 0.021668
107410 1597.8743 1548.8544 1411.974243 1355.757813 0
214950 331.97876 350.3743 328.4424744 324.0109863 0
218620 5576.539 5714.511 5650.027832 5818.570313 0.041596
221100 2574.3164 2713.0063 3260.266113 2742.404053 0.061292
222880 50.29491 138.43388 61.51412582 59.99531174 0.161686
227300 3224.443 3356.0312 3418.429199 3388.108398 0.048306
230410 5307.4634 5733.2676 5856.409668 6040.648926 0.121375
231430 441.79767 377.2723 581.7993774 713.0888062 0.470932
232050 5.8684874 140.72949 6.452753067 6.638870239 0.116041
236390 4528.812 4847.2886 4787.857422 4680.506348 0.03241
236850 1197.9114 1308.7864 1169.570313 1191.723145 0.018589
242760 5888.873 6968.5566 6002.225586 6160.327148 0.044065
244210 657.4606 672.32324 676.5147095 658.0761108 0.000935
250900 895.2101 886.4052 1012.618652 892.572937 0.00691
251570 4304.9175 4886.079 4088.169922 4352.344727 0.060697
252950 1971.4385 1928.533 1971.546509 2054.135742 0.061146
255710 2154.8572 2082.0603 2231.175781 2078.98584 0
270880 789.74634 748.85596 746.2268677 760.31073 0.018524
271590 9292.364 9546.938 10758.82422 9438.995117 0.015535
275850 2671.1484 3121.9731 3179.639404 3118.741699 0.143517
281990 2094.3948 2404.5767 3269.309326 2315.452881 0.095471
284160 929.92413 903.6123 956.4987183 908.8114014 0.005721
289070 4861.033 5243.972 4706.715332 4633.394531 0
291550 1339.3384 1323.4893 1290.662231 1324.343018 0.025432
292030 3181.2307 3723.4417 3607.125732 3500.928223 0.091318
294100 2123.2292 2150.6 1990.630981 2074.705322 0.040524
304930 5698.3926 5597.64 5430.058105 5824.242188 0.06768
306130 1727.7567 1927.9446 1787.971802 1765.812744 0.021552
322170 987.1338 878.4059 902.7176514 910.0273438 0.034748
322330 4585.129 5155.708 4900.762207 4916.681152 0.067434
346110 6389.454 7178.1084 7004.43457 6871.126953 0.070101
359550 5251.984 5247.6694 5184.080566 5156.916016 0
364360 78.73614 182.35359 89.08701324 88.43521118 0.109674
365590 158.19469 229.93842 173.8761902 180.6734314 0.124416
374320 557.42993 630.0981 659.1234131 655.4638672 0.149564
377160 1733.9108 1486.6573 1814.430298 1725.577515 0.138458
381210 5498.251 5568.4336 5691.748047 5964.625488 0.07819
386360 1135.4182 1231.7946 1126.588989 1125.822021 0
394360 2788.8633 2674.6377 2861.023682 2725.428223 0.018636
413150 3061.9893 3204.8018 2831.16748 2713.414551 0
Table 9: Cont. Full result on GAUD dataset in de-normalized MAE. The best result is shown in green shaded bold font. The ones with performance boost over 10% is marked in red.
game_id IATSF_pretrain
IATSF_pretrain
_Gnorm
IATSF PatchTST IMP/%
427520 923.96985 772.3062 803.0150146 775.4301758 0.004029
457140 1159.0262 1217.2543 1133.615601 1107.060547 0
489830 2071.2698 2067.1377 1855.288574 1809.03479 0
493520 270.12488 337.2064 311.7146606 314.7081604 0.141665
513710 1662.712 1732.691 2458.639648 2096.899658 0.207062
526870 1714.113 1704.2557 2002.178223 2031.111084 0.160924
529340 2037.8783 3383.1628 4038.390381 3701.234863 0.449406
548430 3215.861 3539.4717 3559.275391 3622.380615 0.112224
552500 2653.9133 2886.299 3560.82251 4062.790527 0.346776
552990 5025.465 5687.0366 3085.568848 5004.588379 0.383452
578080 21942.736 20576.87 24925.38086 21725.19727 0.052857
582010 2820.5447 3062.556 3161.88623 3088.219727 0.086676
582660 1543.2185 1686.1365 1597.595703 1613.270874 0.043423
646570 1219.2263 1304.4564 1372.947632 1420.136963 0.141473
648800 1754.793 2505.1694 2167.084717 2164.018311 0.189104
739630 3801.0945 4393.0825 4291.587891 4325.070313 0.121149
761890 1032.8406 1291.7933 1046.670288 1021.805847 0
814380 1362.8702 1614.8129 1658.933594 1566.417725 0.129945
892970 3217.5664 2532.8167 3313.418701 3676.412598 0.311063
960090 1899.8359 2076.6377 2291.049316 2292.431396 0.171257
1085660 16838.436 18822.541 18422.99219 18631.20313 0.096224
1091500 13698.44 14906.672 15611.44824 16007.56543 0.144252
1172470 33071.402 39898.113 37495.97266 37135.38672 0.109437
1172620 2595.5396 3012.9434 2990.058838 3043.287109 0.147126
1222670 3100.3125 2550.8154 3211.132813 3179.268799 0.197672
1238810 1900.4874 1945.948 2186.150391 2113.537598 0.100803
1238840 1193.3369 1223.7898 1373.025757 1397.660156 0.14619
1293830 1417.9685 1575.0452 1349.839478 1393.606567 0.031406
1326470 1735.4362 1926.2095 2924.927979 2730.827881 0.364502
1361210 4610.036 4732.349 9080.219727 6853.158203 0.327312
1454400 809.03754 769.7717 941.6995239 1514.508057 0.491735
1623660 576.48615 741.7087 850.6308594 978.7790527 0.411015
1665460 1282.9064 1174.2739 1324.958374 1345.970093 0.127563
1677740 1610.0262 1459.8435 1413.402588 1450.460693 0.025549
1811260 4966.4424 9192.259 8108.043457 7732.84668 0.357747
1868140 1495.137 1424.7719 12159.14551 9878.760742 0.855774
1919590 1274.8295 2378.603 6667.467773 1967.391602 0.35202
1938090 10918.149 10952.329 14469.45508 14213.95605 0.231871
1948980 534.38995 725.96063 1041.738037 1054.809937 0.493378
Best_count 53 17 9 10 0.126324
Appendix N Error Bar & Critical Difference Diagram
We run the experiments on Toy and Electricity for five times with different randomly chosen random seeds. And Weather-Medium for three times because of the large amount of data can result in very long training time on our devices. We report the mean and standard deviation as follows with comparison with FITS, the most stable model.
As Tab. 10 indicate, FIATS shows stable performance across the benchmark. Even with extreme condition, it still maintains superior performance. It worth note that, we thought the relative large variance on weather dataset is caused by the different combination of the text description. But the FITS also shows large variance on this dataset which indicate it is hard to converge on this dataset.
We generate the critical difference plot on our result of four datasets (toy, Electricity, Weather-Medium, Weather-Large) with the default alpha as 0.05 as shown in Fig. 12. FIATS’s placement at the top of the critical difference plot, without intersecting with other lines, demonstrates its consistent and superior performance in terms of MSE compared to the other models. It indicates that with the help of external textual information, FIATS can handle complicated datasets.
Figure 12: The Critical Difference Plot on the FIATS and other baselines with alpha=0.05.
Table 10: The mean and std of FIATS on the three dataset in metrics of MSE.
Datasets FIATS FITS
Toy 0.027±0.001 0.883±0.000
Electricity 0.193±0.004 0.203±0.001
Weather-Medium 0.281±0.008 0.430±0.011
Appendix O Experiment on Time-MMD
Initially, we planned to include the Time-MMD dataset as a real-world scenario to benchmark our method. However, upon evaluation, we found that this dataset suffers from significant flaws and poor organization. As a result, we have decided to include it in the appendix as supplementary material and highlight some of its issues.
O.1 Results on Time-MMD
Despite the dataset’s limitations, our FIATS demonstrates state-of-the-art performance on the Time-MMD dataset, as shown in Tab. 11. In several subsets, FIATS achieves a performance improvement exceeding $50\%$, showcasing its robustness and effectiveness even when applied to a flawed dataset.
Table 11: Results on TimeMMD.
Agriculture Climate
FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG
0.07 0.09 0.17 0.24 1.02 1.24
0.11 0.12 0.19 0.34 1.02 1.24
0.16 0.15 0.24 0.43 1.03 1.24
0.17 0.18 0.29 0.51 1.03 1.24
Economy Traffic
FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG
0.14 0.10 0.35 0.14 0.188 0.24
0.16 0.11 0.4 0.15 0.19 0.25
0.17 0.14 0.41 0.17 0.19 0.24
0.2 0.14 0.37 0.18 0.24 0.29
Socialgood Security
FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG
0.66 0.82 0.87 68.51 112.76 118.49
0.75 0.93 0.99 85.81 115.33 119.09
0.78 1.01 1.07 89.26 117.19 121.08
0.86 1.05 1.13 92.89 118.03 123
Energy Health
FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG
0.11 0.14 0.16 1.38 1.12 1.55
0.21 0.24 0.27 1.82 1.4 1.88
0.3 0.32 0.35 2.01 1.48 1.91
0.36 0.44 0.46 3.04 1.53 1.97
Environment
FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG
0.32 0.32 0.35
0.34 0.35 0.38
0.35 0.37 0.47
0.37 0.41 0.4
O.2 Poor Data Quality of Time-MMD
Unbalanced Dataset.
Many subsets of the Time-MMD dataset span extended periods to collect as many valid numerical data points as possible, with some dating back to the 1980s. However, textual information from earlier periods is largely absent, resulting in a lack of corresponding records for these time intervals. In contrast, recent years have seen a surge in text articles available online. This imbalance creates challenges for the model: during training, it cannot effectively learn correlations between text and time series due to sparse or missing text data, while during inference, the model is inundated with abundant textual information. As a result, the dataset becomes inherently unbalanced.
Meaningless Placeholder Text
Moreover, some entries in the dataset include placeholder text generated by large language models, indicating their inability to produce meaningful output due to insufficient input data. For instance, in the "Agriculture_search.csv" file, over $80\%$ of the entries consist of statements like: "Since there is no relevant information, I am unable to provide any objective facts, insights, analysis, or predictions about the United States broiler market. This search result is not relevant to United States Retail Broiler or Retail Chicken. It appears to be an advertisement for a perfume and has no connection to the topic."
Similarly, in the "Economy_search.csv", other entries state: "After reviewing the search results, I found that most of the information is not relevant to making predictions about the Economy. However, I was able to extract some useful information, which I have summarized below: NA." While these entries appear to be valid text data, they offer no meaningful or actionable information, further compounding the issue of data imbalance.
These meaningless information are all over the whole dataset, making the text validity of this dataset doubtful.
Information Leakage.
Another significant issue with the dataset is information leakage. The dataset creators used large language models to process reports or search results and generate "fact" and "prediction" entries. However, in some cases, the reports or search results directly contain the actual values to be predicted, leading to severe information leakage.
For example, in the "Agriculture_report.csv" file, the "fact" entry for the date 2019-02-04 states: "The National Composite Weighted Average for 1/31/19 is 92.04 compared to 94.22 a week earlier, and 91.66 a year ago." This directly provides the target value to be predicted. Such instances of information leakage are pervasive throughout the dataset and significantly compromise the reliability of the results derived from it.
Appendix P IATSF Benchmark Datasets
We designed the IATSF benchmark to include four datasets of varying complexity, each tailored to evaluate specific aspects of model performance. Together, these datasets form a progression from simple, interpretable scenarios to challenging, real-world applications, providing a comprehensive evaluation framework for text-guided time series forecasting models.
1. Toy Dataset The Toy dataset is intentionally designed with simple and straightforward patterns, making it easy to analyze and interpret. However, the dataset includes sudden changes in patterns that are impossible to predict without text guidance. This ensures the model’s ability to adhere to textual cues is effectively tested in a controlled environment. It serves as a foundation for validating whether the model can extract and use textual guidance to forecast time series.
2. Electricity Dataset The Electricity dataset introduces real-world data with common textual features like day of the week or public holidays. While the textual information is relatively simple, it tests the model’s ability to utilize such structured cues for forecasting. Additionally, as a widely-used off-the-shelf dataset, it allows for easy comparison with existing methods, providing a baseline for evaluating IATSF’s performance.
3. Atmospheric Physics Dataset The Atmospheric Physics dataset represents a semi-controlled environment designed to rigorously test the model’s ability to learn causal relationships between text and time series patterns. It also evaluates the model’s text-guided channel independence and generalizability. By simulating a scenario where text and time series data are strongly correlated, this dataset bridges the gap between controlled tests and more complex real-world challenges.
4. GAUD Dataset The GAUD dataset is a fully real-world dataset that tests IATSF in a practical industrial context. Its patterns are noisy and random, making it highly challenging. This dataset showcases IATSF’s ability to perform well in realistic scenarios.
Comprehensive Benchmark Objectives
The IATSF benchmark is designed to address multiple objectives:
  • Interpretability and Validation: The simpler Toy and Electricity datasets help researchers validate their models and understand their behavior in controlled environments.
  • Performance Testing in Complex Scenarios: The Atmospheric Physics and GAUD datasets challenge the models in semi-controlled and real-world settings, ensuring they are robust and capable of handling practical applications.
This benchmark is not merely a ranking tool but a framework to help researchers analyze and improve their models’ behaviors across varying levels of complexity. We see this as a starting point for the community and hope it will inspire researchers to contribute additional datasets, further expanding and enriching the IATSF benchmark for future advancements in this field.
P.1 Metadata for Datasets
We show the metadata for IATSF Datasets in Table 12.
Table 12: Datasets Metadata
Dataset Length Time span TS Sampling Rate # of Channels # of Dynamic News each step Textual update rate Notes
Toy 300,000 N/A N/A 1 1$\sim$3 Every Step Sinusoidal wave with a single channel
Electrical Utility 26,304 2011-01-01 to 2015-12-31 1 hour 321 1$\sim$3 Daily Just the Electricity Dataset
Atmospheric Physics 525,600 2014-01-01 to 2023-12-31 10 minutes 21 7 Every 6 hours Weather data with 21 channels. Three set of textual cues for combination.
GAUD Varies 2005 to 2024 1 Day 1 (each game) Varies Varies Each game has historical data from its prelaunch to 2024.
P.2 Toy Dataset Details
We directly generate this dataset with sinusoidal wave that randomly changes frequency. Before each changing point, we add 10 captions as ’Channel 1 will change to frequency x in y timesteps.’ After each changing point, we add 5 captions as ’Channel 1 will keep steady with frequency of x.’ In other timesteps, we caption it as ’The waveform will go steady.’
We will publish this dataset with CC BY-NC-SA 4.0 licence.
P.3 Electricity-Caption Details
We caption the day of week with the given time stamp. But we somehow find the original time stamp is incorrect. Instead of the year of 2016, it should be collected in year 2012. Without knowing the exact location of this building, we cannot identify the specific public holiday. We then uses channel 319, which shows obvious patterns of workday and holiday as indicator, when the average value lower than a specific value, we caption it with public holiday.
We will publish this dataset with CC BY-NC-SA 4.0 licence.
P.4 Atmospheric Physics Details
P.4.1 Data source
In creating a IATSF dataset, it is advisable to avoid directly generating the description out of the forecasting horizon time series pattern as news messages, as this could lead to information leakage. News messages should instead contain relevant, known information from other sources. Thus, we get the weather time series data from: https://www.bgc-jena.mpg.de/wetter/ and weather report from https://www.timeanddate.com/weather/germany/jena/historic. We will publish this dataset with CC BY-NC-SA 4.0 license since the data source forbids commercial use.
P.4.2 Motivation
The Atmospheric Physics dataset is designed as a semi-controlled environment to rigorously test the model’s ability to learn causal relationships between text and time series, as well as its text-guided channel independence and generalizability.
Such scenarios are commonly encountered in industrial applications, where correlated text and time series data often coexist. However, obtaining and releasing industrial datasets is challenging due to intellectual property restrictions. To address this, we chose the weather system—a widely available, well-understood, and publicly accessible domain—to simulate these scenarios.
As an off-the-shelf IATSF dataset, the Atmospheric Physics dataset provides a benchmark for evaluating the model’s capacity to learn causal relationships between text and time series patterns, offering a practical and accessible alternative for research and experimentation.
P.4.3 Statistical detail of the Time Series
For better understanding of the statistical distribution of Weather dataset, we plot the histogram of all 21 channels in Fig. 13.
Figure 13: Histogram of all 21 channels. It shows all the channels have unique value distribution, making a model hard to generalize on all channels without knowing related information.
Table 13: Example caption of the Atmospheric Physics dataset.
Topic Example
Month & Time of the Day It’s the early morning of a day in January.
Overall Weather The current weather is clear.
Weather Trend in next 6h The weather is expected to remain clear.
Temperature Trend in next 6h The temperature is showing a mild drop.
Wind Speed & Direction There is Light Breeze from NNW.
Atmosphere Pressure Level The atmospheric shows Average Pressure.
Humidity Level The air is very humid.
P.4.4 Use of Large Language Models for Preprocessing the Weather Dataset
We would like to clarify that the use of Large Language Models (LLMs) in this work is strictly limited to the preprocessing and creation of the Atmospheric Physics dataset. LLMs are not part of our model or method, nor do they contribute to the training or inference process of FIATS. The Atmospheric Physics dataset is intended to serve as an off-the-shelf, text-time synchronized benchmark dataset with raw text and pre-embedded text embeddings as optional inputs.
The primary reason for using LLMs to preprocess this dataset is to generate diverse and correlated textual descriptions, ensuring a richer corpus for training and evaluation. By incorporating varied expressions, we enable the model to generalize to different textual forms while aligning the semantic meaning of text with time series patterns. For instance, the descriptions “The morning will be sunny, but clouds will increase in the afternoon with a chance of light rain” and “The day starts with clear skies, gradually turning cloudy with some rain in the afternoon” carry the same semantic information but differ in expression. This diversity enhances the robustness of the benchmark and validates the model’s generalization capabilities.
Additionally, the raw data source for this dataset often includes general weather reports in text form, accompanied by coarse numerical updates every six hours. While numerical values such as {High_Temp: 25, Low_Temp: 20, Temp_Trend: slightly increasing, Wind_Speed: 5, Wind_Direction: East} are available, they lack the precision required for reliable exogenous variables. Moreover, the raw text contains rich semantic details—such as qualitative weather descriptions—that cannot be effectively captured using numerical values or one-hot encoding. Using text embeddings allows the model to leverage both semantic and numerical information more effectively.
In summary, the LLM preprocessing step is solely for dataset preparation and corpus diversity, ensuring that the Atmospheric Physics dataset is suitable for evaluating text-guided time series forecasting models. Our method does not rely on any LLM capabilities, and the inclusion of LLM-generated text is not a necessary step for IATSF or any similar model. We will include raw data samples in the final paper to provide greater clarity and avoid any misunderstandings.
P.4.5 Channel Details
The meaning of each channel are as follows. The original weather dataset only contains the abbreviation for each channel, to further enrich the semantic for accurate information, we add a line of explanation after it as the channel description.
  • p (mbar): Atmospheric pressure measured in millibars. It indicates the weight of the air above the point of measurement.
  • T (degC): Temperature at the point of observation, measured in degrees Celsius.
  • Tpot (K): Potential temperature, given in Kelvin. This is the temperature that a parcel of air would have if it were brought adiabatically to a standard reference pressure, often used to compare temperatures at different pressures in a thermodynamically consistent way.
  • Tdew (degC): Dew point temperature in degrees Celsius. It’s the temperature to which air must be cooled, at constant pressure and water vapor content, for saturation to occur. A lower dew point means dryer air.
  • rh (%): Relative humidity, expressed as a percentage. It measures the amount of moisture in the air relative to the maximum amount of moisture the air can hold at that temperature.
  • VPmax (mbar): Maximum vapor pressure, in millibars. It represents the maximum amount of moisture that the air can hold at a given temperature.
  • VPact (mbar): Actual vapor pressure, in millibars. It’s the current amount of water vapor present in the air.
  • VPdef (mbar): Vapor pressure deficit, in millibars. The difference between the maximum vapor pressure and the actual vapor pressure; it indicates how much more moisture the air can hold before saturation.
  • sh (g/kg): Specific humidity, the mass of water vapor in a given mass of air, including the water vapor. It’s measured in grams of water vapor per kilogram of air.
  • H2OC (mmol/mol): Water vapor concentration, expressed in millimoles of water per mole of air. It’s another way to quantify the amount of moisture in the air.
  • rho (g/m³): Air density, measured in grams per cubic meter. It indicates the mass of air in a given volume and varies with temperature, pressure, and moisture content.
  • wv (m/s): Wind velocity, the speed of the wind measured in meters per second.
  • max. wv (m/s): Maximum wind velocity observed in the given time period, measured in meters per second.
  • wd (deg): Wind direction, in degrees from true north. This indicates the direction from which the wind is coming.
  • rain (mm): Rainfall amount, measured in millimeters. It indicates how much rain has fallen during the observation period.
  • raining (s): Duration of rainfall, measured in seconds. It specifies how long it has rained during the observation period.
  • SWDR (W/m²): Shortwave Downward Radiation, the amount of solar radiation reaching the ground, measured in watts per square meter.
  • PAR (umol/m2̂/s): Photosynthetically Active Radiation, the amount of light available for photosynthesis, measured in micromoles of photons per square meter per second.
  • max. PAR (umol/m2̂/s): Maximum Photosynthetically Active Radiation observed in the given time period, indicating the peak light availability for photosynthesis.
  • Tlog (degC): Likely a logged temperature measurement in degrees Celsius. It could be a specific type of temperature measurement or recording method used in the dataset.
  • CO2 (ppm): Carbon dioxide concentration in the air, measured in parts per million. It’s a key greenhouse gas and indicator of air quality.
P.4.6 Visualization of the Test Sample
We show a segment of test sample along with the dynamic news timeline in Fig. 14. The news messages are sparse and vague and not directly correlated to some of the channels. These text are passed to the model as text embeddings and aligned with time series on time domain. Thus, the model can extract causal relationship to guide each channel to perform accurate prediction even though they have distinguished distribution.
Figure 14: An visualization of test sample with all the corresponding dynamic news. Atmospheric Physics dataset have dynamic weather report update every 6 hours. As we demonstrate a case of predicting 48 hours. Note that the embedding of these sentences are fed to the FIATS along with the look-back window time series as input. We highlight some of the words that may make impact on the forecasting result.
The following sections give detailed performance and visualization across all the channels.
P.5 GAUD Details
GAUD datasets contains 89 subdataset. Each subdataset correspond to the active user time series of one specific game along with text information. Each subdataset contains one basic information as the channel description includes game title, genera and developer also with the update log includes release date, update type, update title, and article body. We will release the pre embeddings of these text information to avoid violating intellectual property constrains.
Appendix Q Implementation Details and Hyper-Parameters
We train our model on single NVIDIA A800 GPU.
For electricity dataset, we directly report the result from the original paper. For weather dataset, we uses the exact set of hyper-parameter for the original weather datasets provided by each baseline model.
In most of the experiments, we simply use a patch length of 6 and stride of 3. For Toy dataset, we use patch length of 16 and stride of 8.
We follow the previous works, split all the dataset by 7:1:2 for training, validation and testing.
Except the performance on the Atmospheric Physics Dataset, all other experiments are ran on the MiniLM Embedding. We selected MiniLM as the embedding model because it achieves results comparable to OpenAI embeddings while producing smaller embeddings (384 dimensions for MiniLM versus 512 for OpenAI). This reduced embedding size speeds up training, particularly for ablation studies, making it more practical for our experiments.
Further detailed hyperparameter settings are provided in the training scripts in our codebase. We did not perform comprehensive hyper-parameter tuning because of the constraint of compute power. Thus, we may report a sub-optimal result of FIATS.